### v2

# Intuitionism with Truth Tables: A Decision Procedure for IPL Based on RNmatrices

Abstract
This paper provides a new decision procedure for intuitionistic propositional logic (IPL) based on a 3-valued restricted non-deterministic matrix and proves its soundness and completeness.
1 Introduction
A famous result, proved first by K. Gödel in $1932$, says that it is impossible to characterize intuitionistic propositional logic (IPL) via a finite-valued (deterministic) matrix [1]. In other words, Gödel proved that there is no effective decision procedure based on finite-valued ordinary truth-tables for IPL.
Later, in $1940$, a maybe less famous but equally important result was proved by Dugundji, who adapted Gödel’s argument to establish that the same is also true about the family of modal logics between S1 and S5 [1].
To overcome Dugundji’s result, J. Kearns [6] introduced a $4$-valued Nmatrix (non-deterministic matrix) for modal logics KT, S4, and S5 and restricted the set of valuations to a subset called “level valuations”. He proved that this restricted Nmatrix (hereafter called RNmatrix [3]) constitutes sound and complete semantics for these modal logics.
Unfortunately, in principle, Kearns’s level valuations fail to provide an effective decision procedure for any of the modal logics. More than that, until recently, it was unknown whether obtaining such a decision procedure from level valuations would be possible.
(…) the interesting problem of determining whether or not a decision procedure can be obtained by means of level valuations remains open. [2]
L. Grätz recently solved this problem for some modal systems [4]. He refined Kearn’s original RNmatrix to obtain a decidable $3$-valued RNmatrix for modal logics KT and S4. He noticed that, with an appropriate notion of partial valuation for level semantics (i.e., preserving analyticity and co-analyticity), the truth tables could be obtained by removing any spurious row eventually produced in the valuation process.
In this paper, we will try to overcome the original problem posed by Gödel on the characterizability of IPL via finite-valued matrix. We will propose a decidable 3-valued RNmatrix for IPL combining Grätz’s RNmatrix for S4 and Gödel-McKinsey-Tarski’s box translation.
2 Nmatrices and RNmatrices
Along this paper, all the propositional languages will be defined over a given denumerable set $\mathcal{P}=\{p_{0},p_{1},p_{2},\ldots\}$ of propositional variables.
Definition 2.1 (Nmatrix).
Let $\Theta$ be a signature and $For(\Theta)$ the propositional language over signature $\Theta$ generated by the set $\mathcal{P}$ of propositional variables. An Nmatrix for $\mathcal{L}$ over $\Theta$ is a tuple $\mathcal{M}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$ where
  1. 1.
    $\mathcal{V}$ is the (non-empty) set of truth values;
  2. 2.
    $\mathcal{D}\subset\mathcal{V}$ is the (non-empty) set of designated values;
  3. 3.
    $\mathcal{O}$ is a function assigning, to each $n$-ary connective ##\## of $\Theta$, a (non-deterministic) truth function $\mathcal{O}(\#):\mathcal{V}^{m}\rightarrow(\wp(\mathcal{V}))\backslash\emptyset$.
A valuation over an Nmatrix $\mathcal{M}$ is a function $v:For(\Theta)\to\mathcal{V}$ such that, for every $n$-ary connective #normal-#\## of $\Theta$ and for every $\varphi_{1},\ldots,\varphi_{n}\in For(\Theta)$, $v(\#(\varphi_{1},\ldots,\varphi_{n}))\in\mathcal{O}(\#)(v(\varphi_{1}),\ldots,v(\varphi_{n}))$. The set of valuations over $\mathcal{M}$ is denoted by $Val(\mathcal{M})$.
Definition 2.2 (RNmatrix).
Let $\Theta$ be a signature. An RNmatrix [3] over $\Theta$ is a pair $\mathcal{R}=\langle\mathcal{M},\mathcal{F}\rangle$ where $\mathcal{M}$ is an Nmatrix over $\Theta$ and $\mathcal{F}\subseteq Val(\mathcal{M})$.
Definition 2.3 (Consequence relation in an (R)Nmatrix).
Let $\mathcal{M}$ ($\mathcal{R}$, resp.) be an Nmatrix (an RNmatrix, resp.) over $\Theta$, and $\Gamma\cup\{\varphi\}\subseteq For(\Theta)$. We say that $\varphi$ is a semantical consequence of $\Gamma$ w.r.t. $\mathcal{M}$ (w.r.t. $\mathcal{R}$, resp.), denoted by $\Gamma\models_{\mathcal{M}}\varphi$ (denoted by $\Gamma\models_{\mathcal{R}}\varphi$, resp.) if, for every $v\in Val(\mathcal{M})$ (for every $v\in\mathcal{F}$, resp.): $v(\varphi)\in\mathcal{D}$ whenever $v(\gamma)\in\mathcal{D}$ for every $\gamma\in\Gamma$. We write $\models_{\mathcal{M}}\varphi$ and $\models_{\mathcal{R}}\varphi$ instead of $\emptyset\models_{\mathcal{M}}\varphi$ and $\emptyset\models_{\mathcal{R}}\varphi$.
$\neg^{\mathcal{A}_{S4}}$
0 $\{1,2\}$
1 $\{0\}$
2 $\{0\}$
$\Box^{\mathcal{A}_{S4}}$
0 $\{0\}$
1 $\{0\}$
2 $\{2\}$
$\rightarrow^{\mathcal{A}_{S4}}$ 0 1 2
0 $\{1,2\}$$\{1,2\}$$\{1,2\}$
1 $\{0\}$$\{1,2\}$$\{1,2\}$
2 $\{0\}$$\{1\}$$\{1,2\}$
$\lor^{\mathcal{A}_{S4}}$ 0 1 2
0 $\{0\}$$\{1,2\}$$\{1,2\}$
1 $\{1,2\}$$\{1,2\}$$\{1,2\}$
2 $\{1,2\}$$\{1,2\}$$\{1,2\}$
$\land^{\mathcal{A}_{S4}}$ 0 1 2
0 $\{0\}$$\{0\}$$\{0\}$
1 $\{0\}$$\{1,2\}$$\{1,2\}$
2 $\{0\}$$\{1,2\}$$\{1,2\}$
Table 1: Tables for $\mathcal{M}_{S4}$.
Example 2.4 ($\mathcal{M}_{S4}$).
Let $\Sigma_{b}=\{\{\neg,\Box\},\{\rightarrow\}\}$ and Σ={{¬,},{,,}}normal-$\Sigma=\{\{\neg,\Box\},\{\rightarrow,\lor,\land\}\}$ be, respectively, the basic signature and the signature for modal logic S4. Given the set $\mathcal{P}$ of propositional variables, let $For(\Sigma_{b})$ and $For(\Sigma)$ be the corresponding propositional languages. Grätz’s Nmatrix over $\Sigma_{b}$ is the tuple $\mathcal{M}_{S4}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$, where $\mathcal{V}=\{0,1,2\}$, $\mathcal{D}=\{1,2\}$, $\mathcal{O}=\{\neg^{\mathcal{A}_{S4}},\Box^{\mathcal{A}_{S4}},\rightarrow^{\mathcal{A}_{S4}}\}$ and each $c\in\mathcal{O}$ is defined in Table 1.
$p$$\Box p$$\Box\Box p$$\Box p\rightarrow\Box\Box p$000121001222212
(a) Valuations in $\mathcal{M}_{S4}$ for $\Box p\rightarrow\Box\Box p$
Figure 1: Each tree level corresponds to the set of values resulting from the multioperation applied to the respective subformula (at left) in that tree. At the top are the values that the atomic formula can receive, and in the leaves are the values of the evaluated formula.
By setting $\varphi\lor\psi:=\neg\varphi\to\psi$ and $\varphi\land\psi:=\neg(\varphi\to\neg\psi)$, respectively, the multioperators $\lor^{\mathcal{A}_{S4}}$ and $\land^{\mathcal{A}_{S4}}$ are defined as in Table 1.
Valuations in an Nmatrix are typically represented as tables. However, if one focuses more on the computational aspects of these valuations, adopting the more flexible data structure of binary trees can be advantageous. In Figure 1, we show how to compute the valuation of two examples.
3 Kearns’s level valuations
Regarding semantics for modal logic, Kripke-style semantics are the usual choice. However, there are other attractive alternatives. In the 1980s, Kearns proposed an RNmatrix-based semantics for modal logics KT, S4 and S5. He introduced the level valuations as the set $\mathcal{F}$ of valuations to be used.
[Kearn’s account] is simpler than [Kripke’s account] in virtue of having dispensed with possible worlds and their relations. I also think that my account is philosophically preferable to the standard account for having done this. For I do not think there are such things as possible worlds, or even that they constitute a useful fiction. [6]
We will define the level valuations for the modal logic S4.
Figure 2: Assume that $Val(\mathcal{M}_{S4})=\{v_{0},v_{1},\ldots,v_{m},v_{n},v_{k},\ldots\}$. We remove a valuation $v$ in a level $\mathcal{L}_{i+1}$ iff there is some wf $\alpha$ such that $v(\alpha)=1$ and $w(\alpha)\in\mathcal{D}$ for every $w\in\mathcal{L}_{i}$.
Definition 3.1.
(Level valuations for S4)
Assume that $Val(\mathcal{M}_{S4})$ is the set of valuations over $\mathcal{M}_{S4}$. We define the level $\mathcal{L}^{S4}_{k}$, where $k\in\mathbb{N}$, as follows:
  1. i)
    $\mathcal{L}_{0}^{S4}=Val(\mathcal{M}_{S4})$;
  2. ii)
    $\mathcal{L}^{S4}_{k+1}=\{v\in\mathcal{L}^{S4}_{k}\mid\forall\alpha\in For(\Sigma_{b}),\mathcal{L}^{S4}_{k}(\alpha)\subseteq\mathcal{D}\Rightarrow v(\alpha)=2\}$
where $\mathcal{L}^{S4}_{k}(\alpha)=\{v(\alpha)\mid v\in\mathcal{L}^{S4}_{k}\}$.
Definition 3.2.
The set of level valuations in $\mathcal{M}_{S4}$ is defined as the intersection of every level valuation:
$\mathcal{L}_{S4}=\cap^{\infty}_{k\geq 0}\mathcal{L}^{S4}_{k}$
Definition 3.3 (Grätz’s RNmatrix for S4).
Let $\mathcal{M}_{S4}$ be the nmatrix defined in Definition 2.4. The RNmatrix for S4 is the pair $\mathcal{R}(\mathcal{M}_{S4})=\langle\mathcal{M}_{S4},\mathcal{L}_{S4}\rangle$.
Let $\vdash_{S4}^{b}$ be the consequence relation of the standard Hilbert calculus $\mathcal{H}^{b}_{S4}$ for S4 over $\Sigma_{b}$. As usual, $\Gamma\vdash_{S4}^{b}\varphi$ iff either $\varphi$ is a theorem of $\mathcal{H}^{b}_{S4}$ or there exists a non-empty set $\{\beta_{1},\ldots,\beta_{n}\}\subseteq\Gamma$ such that $\beta_{1}\to(\beta_{2}\to(\ldots\to(\beta_{n}\to\varphi)\ldots))$ is a theorem of $\mathcal{H}^{b}_{S4}$.
Theorem 3.4 ([4]).
For every $\Gamma$ and $\varphi$: $\Gamma\vdash_{S4}^{b}\varphi$ iff $\Gamma\models_{\mathcal{R}(\mathcal{M}_{S4})}\varphi$.
3.1 Grätz’s decision procedure for S4
While the soundness and completeness of level valuations for S4 (with the rule of necessitation) were already known, it was only recently that an effective decision procedure was discovered.
Before that, J. V. Ivlev seems to have suggested the following approach [4]: remove every partial valuation $v_{p}$ whenever (1) $v_{p}(\Box A)\not\in D$ and (2) $w_{p}(A)\in D$ for every partial valuation $w_{p}$. The reason for this method is very intuitive if we consider the definition of level valuations, and indeed, it works for $2$-valued semantics, as well as for some examples in $\mathcal{M}_{S4}$, such as $\Box\alpha$ where $\alpha$ is any propositional tautology. However, this criterion is not sufficient for $\mathcal{M}_{S4}$ because it leaves behind invalid partial level valuations (consider the valid formula $\Box p\rightarrow\Box\neg\neg p$ in $\mathcal{M}_{S4}$).
In [4], Grätz identified the need to find a suitable method by refining the concept of partial valuation. To address this, he suggests a new definition for partial level valuations:
Definition 3.5 (Partial valuation in $\mathcal{M}_{S4}$).
A partial valuation in $\mathcal{M}_{S4}$ is a function $\tilde{v}_{p}:\Lambda\rightarrow\mathcal{V}$ such that $\Lambda\subseteq For(\Sigma_{b})$ is closed under subformulas [^4] That is: if $\alpha\in\Lambda$ and $\beta$ is a subformula of $\alpha$ then $\beta\in\Lambda$. and, for every $\alpha,\beta\in\Lambda$:
[…]
  • if $\alpha\to\beta\in\Lambda$ then $\tilde{v}_{p}(\alpha\to\beta)\in{\to}^{\mathcal{A}_{S4}}(\tilde{v}_{p}(\alpha),\tilde{v}_{p}(\beta))$.
Definition 3.6 (Grätz’s partial level valuation).
A partial valuation $v_{p}:\Lambda\rightarrow\{0,1,2\}$ in $\mathcal{M}_{S4}$ is a partial level valuation in $\mathcal{R}(\mathcal{M}_{S4})$ ($v_{p}\in PLV$) iff
  1. 1.
    $\forall\alpha\in\Lambda$ such that $v_{p}(\alpha)=1$, there exists $w_{p}\in PLV$ such that $w_{p}(\alpha)=0$ and, $\forall\beta\in\Lambda$, $w_{p}(\beta)=2$ whenever $v_{p}(\beta)=2$.
Remark 3.7.
In [4, p. 15] it is observed that, being $\Lambda$ finite, the set of $PLV$s with domain $\Lambda$ is well-defined (that is, it is not cyclic), given that the search space is finite: the set of all the partial valuations with domain $\Lambda$ is finite. Moreover, it is proved in [4] that this notion is analytic in $\mathcal{M}_{S4}$ (i.e, every partial level valuation in $\mathcal{M}_{S4}$ can be extended to a level valuation in $\mathcal{M}_{S4}$) and co-analytic (i.e, for any level valuation $v$ in $\mathcal{M}_{S4}$, the restriction of $v$ to a partial level valuation $v_{p}$ defined over a finite set $\Lambda$ closed under subformulas is a partial level valuation).
Theorem 3.8.
$\mathcal{R}(\mathcal{M}_{S4})$ is analytic and co-analytic with PLV.
Using this, one can easily prove that the truth table method is sound and complete, that is,
Theorem 3.9 ([4]).
For every $\varphi$: $\models_{\mathcal{R}(\mathcal{M}_{S4})}\varphi$ if, and only if,for every partial valuation $v_{p}$ in $\mathcal{M}_{S4}$ defined over the set $\Lambda$ of subformulas of $\varphi$: if $v_{p}\in PLV$, then $v_{p}(\varphi)\in\mathcal{D}$.
Corollary 3.9.1 ([4]).
For every $\varphi$: $\vdash_{S4}^{b}\varphi$ if, and only if, for every partial valuation $v_{p}$ in $\mathcal{M}_{S4}$ defined over the set $\Lambda$ of subformulas of $\varphi$: if $v_{p}\in PLV$ then $v_{p}(\alpha)\in\mathcal{D}$.
Proof.
It is an immediate consequence of Theorems 3.9 and 3.4. ∎
Therefore, to construct a truth table, compute every partial valuation and remove any non-PLV. And this algorithmic process is a sound and complete decision procedure for S4. Obviously the soundness and completeness Theorem 3.9 (and so, the decision procedure) can be extended to inferences in S4 from a finite set of premises, taking into account that $\vdash_{S4}^{b}$ (and so $\models_{\mathcal{R}(\mathcal{M}_{S4})}$) satisfies the deduction metatheorem.
3.2 Reducing Grätz’s RNmatrix and algorithm for S4
As suggested by Grätz ([4], p. 17), it is possible to apply the notion of partial level valuation to the tables for the connectives of S4 (in our case, this includes conjunction and disjunction). Then, by eliminating rows $v_{p}$ which are not supported by any other $w_{p}$ as prescribed in Definition 3.6, it is obtained a reduced truth-table for each multioperator of S4 over the full signature $\Sigma$. Along this section we will deal exclusively with signature $\Sigma$ for S4.
[…]
$\neg^{\mathcal{A}_{S4}}$
0 $\{1,2\}$
1 $\{0\}$
2 $\{0\}$
$\Box^{\mathcal{A}_{S4}}$
0 $\{0\}$
1 $\{0\}$
2 $\{2\}$
$\rightarrow^{\mathcal{A}^{\prime}_{S4}}$ 0 1 2
0 $\{1,2\}$$\{1,2\}$$\{2\}$
1 $\{0\}$$\{1,2\}$$\{2\}$
2 $\{0\}$$\{1\}$$\{2\}$
$\lor^{\mathcal{A}^{\prime}_{S4}}$ 0 1 2
0 $\{0\}$$\{1,2\}$$\{2\}$
1 $\{1,2\}$$\{1,2\}$$\{2\}$
2 $\{2\}$$\{2\}$$\{2\}$
$\land^{\mathcal{A}^{\prime}_{S4}}$ 0 1 2
0 $\{0\}$$\{0\}$$\{0\}$
1 $\{0\}$$\{1\}$$\{1\}$
2 $\{0\}$$\{1\}$$\{2\}$
Table 2: Tables for $\mathcal{M}^{\prime}_{S4}$.
Of course this move requires adjusting the soundness and completeness proofs of Grätz to the new RNmatrix defined over signature $\Sigma$. Let $\vdash_{S4}$ be the consequence relation of the standard Hilbert calculus $\mathcal{H}_{S4}$ for S4 over $\Sigma$. The notion $\Gamma\vdash_{S4}\varphi$ of derivations from premises in $\mathcal{H}_{S4}$ is defined, as in the case of $\mathcal{H}^{b}_{S4}$, in terms of theoremhood.
Let $Val(\mathcal{M}^{\prime}_{S4})$ is the set of valuations over $\mathcal{M}^{\prime}_{S4}$. Consider the family of level $\mathcal{L}^{\prime S4}_{k}$ ($k\geq 0$) as in Definition 3.1, but now starting from ${\mathcal{L}^{\prime}}_{0}^{S4}=Val(\mathcal{M}^{\prime}_{S4})$ in item i), and considering formulas $\alpha\in For(\Sigma)$ in item ii). Let $\mathcal{L}^{\prime}_{S4}=\cap^{\infty}_{k\geq 0}\mathcal{L}^{\prime S4}_{k}$ and $\mathcal{R}(\mathcal{M}^{\prime}_{S4})=\langle\mathcal{M}^{\prime}_{S4},\mathcal{L}^{\prime}_{S4}\rangle$.
Remark 3.10.
Let $\mathcal{M}$ be an Nmatrix. By the very definitions it is immediate to see that, if $\mathcal{F}\subseteq\mathcal{F}^{\prime}\subseteq Val(\mathcal{M})$, then the following holds: $\Gamma\models_{\langle\mathcal{M},\mathcal{F}^{\prime}\rangle}\varphi$ implies that $\Gamma\models_{\langle\mathcal{M},\mathcal{F}\rangle}\varphi$.
Theorem 3.11 (Soundness of $\mathcal{H_{S4}}$ w.r.t. $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$).
For every $\Gamma$ and $\varphi$: $\Gamma\vdash_{S4}\varphi$ implies that $\Gamma\models_{\mathcal{R}(\mathcal{M}^{\prime}_{S4})}\varphi$.
Proof.
It coincides with the proof for $\mathcal{R}(\mathcal{M}_{S4})$ found in [4, Theorems 3.4 and 3.5]. The only detail to be checked is that any axiom of $\mathcal{H_{S4}}$ is still valid in the reduced Nmatrix $\mathcal{M}^{\prime}_{S4}$. Details are left to the reader. ∎
Corollary 3.11.1.
If $\vdash_{S4}\varphi$ then $v(\varphi)=2$ for every $v\in\mathcal{L}^{\prime}_{S4}$.
Proof.
Assume that $\vdash_{S4}\varphi$, and let $v\in\mathcal{L}^{\prime}_{S4}$. By Theorem 3.11, $v(\varphi)\in\{1,2\}$. Since $\vdash_{S4}\Box\varphi$ (by necessitation rule) then $v(\Box\varphi)\in\tilde{\Box}v(\varphi)$ such that $v(\Box\varphi)\in\{1,2\}$, by Theorem 3.11. Hence, $v(\varphi)=2$. ∎
For completeness, recall that, given a Tarskian and finitary logic L and a formula $\varphi$, a set $\Delta$ of formulas is said to be $\varphi$-saturated in L if $\Delta\nvdash_{\bf L}\varphi$, but $\Delta,\alpha\vdash_{\bf L}\varphi$, whenever $\alpha\notin\Delta$. It is a well-known result that, if $\Gamma\nvdash_{\bf L}\varphi$, there exists a $\varphi$-saturated set $\Delta$ in L such that $\Gamma\subseteq\Delta$. If $\Delta$ is $\varphi$-saturated in L then it is closed, that is: $\Delta\vdash_{\bf L}\alpha$ iff $\alpha\in\Delta$. In particular, this holds in the logic S4, which is generated by $\mathcal{H_{S4}}$ over $\Sigma$. Since it contains classical logic, any $\varphi$-saturated set $\Delta$ satisfies: $\alpha\vee\beta\in\Delta$ iff $\alpha\in\Delta$ or $\beta\in\Delta$; $\alpha\land\beta\in\Delta$ iff $\alpha\in\Delta$ and $\beta\in\Delta$; and $\alpha\to\beta\in\Delta$ iff either $\alpha\notin\Delta$ or $\beta\in\Delta$. The following result is an adaptation of Lemma 3.7 in [4] to the reduced Nmatrix $\mathcal{M}^{\prime}_{S4}$:
Lemma 3.12.
Let $\Delta\subseteq For(\Sigma)$ be a $\varphi$-saturated set in the logic generated by $\mathcal{H_{S4}}$ (that is, S4). Let $v_{\Delta}:For(\Sigma)\to\mathcal{V}$ be the function defined as follows:
[…]
Then, $v_{\Delta}$ is a valuation over the Nmatrix $\mathcal{M}^{\prime}_{S4}$.
Proof.
[…]
Disjunction: Suppose that $v_{\Delta}(\alpha\vee\beta)=2$. Then, $\Box(\alpha\vee\beta)\in\Delta$ and so $\alpha\vee\beta\in\Delta$, by axiom T: $\Box\psi\to\psi$ and the fact that $\Delta$ is closed under logical inferences in $\mathcal{H_{S4}}$. This implies that $\alpha\in\Delta$ or $\beta\in\Delta$, hence $v_{\Delta}(\alpha)\in\mathcal{D}$ or $v_{\Delta}(\beta)\in\mathcal{D}$. This implies that $v_{\Delta}(\alpha\vee\beta)=2\in\vee^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Suppose now that $v_{\Delta}(\alpha\vee\beta)=1$. Then, $\Box(\alpha\vee\beta)\notin\Delta$ but $\alpha\vee\beta\in\Delta$ and so $\alpha\in\Delta$ or $\beta\in\Delta$. Suppose that $\Box\alpha\in\Delta$ or $\Box\beta\in\Delta$. Then, $\Box\alpha\vee\Box\beta\in\Delta$, and so $\Box(\alpha\vee\beta)\in\Delta$, given that $(\Box\alpha\vee\Box\beta)\to\Box(\alpha\vee\beta)$ is a theorem in $\mathcal{H_{S4}}$. But this is a contradiction, therefore $\Box\alpha\notin\Delta$ and $\Box\beta\notin\Delta$. This means that $v_{\Delta}(\alpha)=1$ and $v_{\Delta}(\beta)\in\{0,1\}$, or vice versa. Then, $v_{\Delta}(\alpha\vee\beta)=1\in\{1,2\}=\vee^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Finally, if $v_{\Delta}(\alpha\vee\beta)=0$ then $\alpha\vee\beta\notin\Delta$, hence $\alpha\notin\Delta$ and $\beta\notin\Delta$. This means that $v_{\Delta}(\alpha)=v_{\Delta}(\beta)=0$ and so $v_{\Delta}(\alpha\vee\beta)=0\in\{0\}=\vee^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
Conjunction: Suppose that $v_{\Delta}(\alpha\land\beta)=2$. Then, $\Box(\alpha\land\beta)\in\Delta$ and so $\Box\alpha\land\Box\beta\in\Delta$, since $\Box(\alpha\land\beta)\to(\Box\alpha\land\Box\beta)$ is a theorem in $\mathcal{H_{S4}}$. This implies that $\Box\alpha\in\Delta$ and $\Box\beta\in\Delta$, hence $v_{\Delta}(\alpha)=v_{\Delta}(\beta)=2$. Thus, $v_{\Delta}(\alpha\land\beta)=2\in\{2\}=\land^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Now, assume that $v_{\Delta}(\alpha\land\beta)=1$. Then, $\Box(\alpha\land\beta)\notin\Delta$ but $\alpha\land\beta\in\Delta$ and so $\alpha\in\Delta$ and $\beta\in\Delta$ but either $\Box\alpha\notin\Delta$ or $\Box\beta\notin\Delta$ (since $(\Box\alpha\land\Box\beta)\to\Box(\alpha\land\beta)$ is a theorem in $\mathcal{H_{S4}}$). That is, $v_{\Delta}(\alpha),v_{\Delta}(\beta)\in\{1,2\}$ and either $v_{\Delta}(\alpha)=1$ or $v_{\Delta}(\beta)=1$. Then, $v_{\Delta}(\alpha\land\beta)=1\in\{1\}=\land^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Finally, if $v_{\Delta}(\alpha\land\beta)=0$ then $\alpha\land\beta\notin\Delta$, hence $\alpha\notin\Delta$ or $\beta\notin\Delta$. This means that $v_{\Delta}(\alpha)=0$ or $v_{\Delta}(\beta)=0$ and so $v_{\Delta}(\alpha\land\beta)=0\in\{0\}=\land^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
Implication: Suppose that $v_{\Delta}(\alpha\to\beta)=2$. Then, $\Box(\alpha\to\beta)\in\Delta$ and so $\alpha\to\beta\in\Delta$, by axiom T. This implies that either $\alpha\notin\Delta$ or $\beta\in\Delta$. If $\alpha\notin\Delta$ then $v_{\Delta}(\alpha)=0$. Thus, $v_{\Delta}(\alpha\to\beta)=2\in{\to}^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. If $\beta\in\Delta$, suppose first that $\Box\alpha\in\Delta$. Then, by axiom K it follows that $\Box\beta\in\Delta$. Hence, $v_{\Delta}(\alpha)=v_{\Delta}(\beta)=2$ and so $v_{\Delta}(\alpha\to\beta)=2\in\{2\}={\to}^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Otherwise, if $\Box\alpha\notin\Delta$ then $v_{\Delta}(\alpha)\neq 2$ and $v_{\Delta}(\beta)\neq 0$. From this, $v_{\Delta}(\alpha\to\beta)=2\in{\to}^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Suppose now that $v_{\Delta}(\alpha\to\beta)=1$. Then, $\Box(\alpha\to\beta)\notin\Delta$ but $\alpha\to\beta\in\Delta$ and so $\alpha\notin\Delta$ or $\beta\in\Delta$. In addition, $\Box\beta\notin\Delta$, since $\Box\beta\to\Box(\alpha\to\beta)$ is a theorem in $\mathcal{H_{S4}}$. If $\alpha\notin\Delta$ then $v_{\Delta}(\alpha)=0$ and $v_{\Delta}(\beta)\in\{0,1\}$. Thus, $v_{\Delta}(\alpha\to\beta)=1\in\{1,2\}={\to}^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Otherwise, if $\beta\in\Delta$ then $v_{\Delta}(\beta)=1$ and so $v_{\Delta}(\alpha\to\beta)=1\in{\to}^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$. Finally, if $v_{\Delta}(\alpha\to\beta)=0$ then $\alpha\to\beta\notin\Delta$, hence $\alpha\in\Delta$ and $\beta\notin\Delta$. This means that $v_{\Delta}(\alpha)\in\{1,2\}$ and $v_{\Delta}(\beta)=0$ and so $v_{\Delta}(\alpha\to\beta)=0\in\{0\}={\to}^{\mathcal{A}^{\prime}_{S4}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
This completes the proof. ∎
Theorem 3.13 (Completeness of $\mathcal{H_{S4}}$ w.r.t. $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$).
For every $\Gamma$ and $\varphi$: $\Gamma\models_{\mathcal{R}(\mathcal{M}^{\prime}_{S4})}\varphi$ implies that $\Gamma\vdash_{S4}\varphi$.
Proof.
It is analogous to the proof for $\mathcal{R}(\mathcal{M}_{S4})$ found in [4, Lemma 3.9 and Theorem 3.10], but now by using our Lemma 3.12 instead of [4, Lemma 3.7]. ∎
Now, Definitions 3.5 and 3.6 can be easily adapted to $\mathcal{M}^{\prime}_{S4}$, obtaining so the set $PLV^{\prime}$ of partial level valuations in $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$ over finite sets of formulas. However, it will be convenient to give a bit more detailed presentation of these sets, by specifying explicitly the domain of each valuation.
Definition 3.14.
Consider the sets
[…]
Definition 3.15 (Partial valuation in $\mathcal{M}^{\prime}_{S4}$).
Let $\Lambda\in CS(\Sigma)$. A partial valuation in $\mathcal{M}^{\prime}_{S4}$ is a function $\tilde{v}_{p}:\Lambda\rightarrow\mathcal{V}$ such that, for every $\alpha,\beta\in\Lambda$:
[…]
Let $PV(\Lambda)$ be the set of partial valuations in $\mathcal{M}^{\prime}_{S4}$ with domain $\Lambda$.
Definition 3.16 (Partial level valuation over $\Lambda$).
Let $\Lambda\in FCS(\Sigma)$. A partial valuation $\tilde{v}_{p}\in PV(\Lambda)$ is a partial level valuation in $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$ over $\Lambda$ iff
  1. $\forall\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=1$, there exists a partial level valuation $\tilde{w}_{p}$ in $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$ over $\Lambda$ such that $\tilde{w}_{p}(\alpha)=0$ and, $\forall\beta\in\Lambda$, $\tilde{w}_{p}(\beta)=2$ whenever $\tilde{v}_{p}(\beta)=2$.
The set of partial level valuations in $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$ over $\Lambda$ will be denoted by $PLV(\Lambda)$.
Remark 3.17.
Given $v,w$ and $\alpha$ let $\mathsf{P}_{\Lambda}(v,w,\alpha)$ iff $w(\alpha)=0$ and, $\forall\beta\in\Lambda$, $w(\beta)=2$ whenever $v(\beta)=2$. If $\Lambda\in FCS(\Sigma)$ then
$\begin{array}[]{lll}PLV(\Lambda)&=&\{\tilde{v}_{p}\in PV(\Lambda)\mid\forall\alpha\in\Lambda\big{(}\tilde{v}_{p}(\alpha)=1\mbox{ implies that }\mathsf{P}_{\Lambda}(\tilde{v}_{p},\tilde{w}_{p},\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $\tilde{w}_{p}\in PLV(\Lambda)$}\big{)}\}.\end{array}$
As observed in Remark 3.7 for $\mathcal{M}_{S4}$, since $\Lambda$ is finite then $PV(\Lambda)$ is finite, hence the definition above is not cyclic.
The next step is to show analyticity and co-analyticity w.r.t. level valuations in $\mathcal{M}^{\prime}_{S4}$. In order to do this it is considered, adapting [4, Definition 3.12], the intermediary notion of partial’ level valuation in $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$ over $\Lambda$.
Definition 3.18 (Partial’ level valuation over $\Lambda$).
Let $\Lambda\in CS(\Sigma)$. A partial valuation $\tilde{v}_{p}\in PV(\Lambda)$ is a partial’ level valuation in $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$ over $\Lambda$ iff
  1. $\forall\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=1$, there exists a level valuation $w$ in $\mathcal{L}^{\prime}_{S4}$ such that $w(\alpha)=0$ and, $\forall\beta\in\Lambda$, $w(\beta)=2$ whenever $\tilde{v}_{p}(\beta)=2$.
The set of partial’ level valuations in $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$ over $\Lambda$ will be denoted by $PLV^{\prime}(\Lambda)$.
Remark 3.19.
Consider the predicate $\mathsf{P}_{\Lambda}(v,w,\alpha)$ introduced in Remark 3.17. Clearly,
$\begin{array}[]{lll}PLV^{\prime}(\Lambda)&=&\{\tilde{v}_{p}\in PV(\Lambda)\mid\forall\alpha\in\Lambda\big{(}\tilde{v}_{p}(\alpha)=1\mbox{ implies that }\mathsf{P}_{\Lambda}(\tilde{v}_{p},w,\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $w\in\mathcal{L}^{\prime}_{S4}$}\big{)}\}.\end{array}$
Lemma 3.20 (Co-analyticity lemma for $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$).
Let $v\in\mathcal{L}^{\prime}_{S4}$ and $\Lambda\in CS(\Sigma)$. Then, the restriction $\tilde{v}_{p}:=v_{|\Lambda}$ of $v$ to the domain $\Lambda$ belongs to $PLV^{\prime}(\Lambda)$.
Proof.
It is an adaptation of the proof of [4, Lemma 3.13]. Suppose, by contradiction, that $\tilde{v}_{p}\notin PLV^{\prime}(\Lambda)$. Let $\Gamma:=\{\beta\in\Lambda\mid\tilde{v}_{p}(\beta)=2\}$. Then, there exists $\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=1$ and, for every $w\in\mathcal{L}^{\prime}_{S4}$, $w(\alpha)=0$ implies that $w(\beta)\neq 2$ for some $\beta\in\Gamma$. Hence, for every $w\in\mathcal{L}^{\prime}_{S4}$, $w(\alpha)\not\in\mathcal{D}$ implies that $w(\Box\beta)\notin\mathcal{D}$ for some $\beta\in\Gamma$. By contraposition, this means that $\{\Box\beta\mid\beta\in\Gamma\}\models_{\mathcal{R}(\mathcal{M}^{\prime}_{S4})}\alpha$. By Theorem 3.13, $\{\Box\beta\mid\beta\in\Gamma\}\vdash_{S4}\alpha$. Observe that $\nvdash_{S4}\alpha$: otherwise, $v(\alpha)=2$, by Corollary  3.11.1. Hence, there exist $\beta_{1},\ldots,\beta_{n}\in\Gamma$ such that $\vdash_{S4}\Box\beta_{1}\to(\Box\beta_{2}\to(\ldots\to(\Box\beta_{n}\to\alpha)\ldots))$. By necessitation rule, $\vdash_{S4}\Box(\Box\beta_{1}\to(\Box\beta_{2}\to(\ldots\to(\Box\beta_{n}\to\alpha)\ldots)))$. By Theorem 3.11, $\models_{\mathcal{R}(\mathcal{M}^{\prime}_{S4})}\Box(\Box\beta_{1}\to(\Box\beta_{2}\to(\ldots\to(\Box\beta_{n}\to\alpha)\ldots)))$. But $v(\Box\beta_{i})=2$ for $1\leq i\leq n$ and $v(\alpha)=1$, hence $v(\Box\beta_{1}\to(\Box\beta_{2}\to(\ldots\to(\Box\beta_{n}\to\alpha)\ldots)))=1$. This means that $v(\Box(\Box\beta_{1}\to(\Box\beta_{2}\to(\ldots\to(\Box\beta_{n}\to\alpha)\ldots))))=0$, a contradiction. From this we infer that $\tilde{v}_{p}\in PLV^{\prime}(\Lambda)$. ∎
Definition 3.21.
[…]
Remark 3.22.
Observe that it is possible to define an enumeration $\alpha_{1},\alpha_{2},\ldots\alpha_{m},\ldots$ (for $m\in\omega^{2}$) of $For(\Sigma)$ such that $\textsf{c}_{1}(\alpha_{i})\leq\textsf{c}_{1}(\alpha_{j})$ if $i\leq j$ and, for every $i$ such that $\textsf{c}_{1}(\alpha_{i})>0$, if $\beta$ is a strict subformula of $\alpha_{i}$ then $\beta=\alpha_{j}$ for some $j<i$. Such an enumeration of $For(\Sigma)$ can be defined as follows: every formula have and index in $\omega^{2}$, which is a denumerable ordinal, such that all the formulas with complexity 00 (which form a denumerable set) are placed first, with indexes in $I_{0}:=\omega$; after this, all the formulas with complexity $1$ (which form a denumerable set) are placed with an index in $I_{1}:=\omega\cdot 2\setminus\omega=\{\omega,\omega+1,\ldots\}$; in general, the formulas with complexity $n$ (which form a denumerable set) have an index in $I_{n}:=\omega\cdot(n+1)\setminus\omega\cdot n=\{\omega\cdot n,\omega\cdot n+1,\ldots\}$.
Lemma 3.23 (Analyticity lemma for $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$).
Let $\Lambda\in CS(\Sigma)$ and $\tilde{v}_{p}\in PLV^{\prime}(\Lambda)$. Then, there exists $v\in\mathcal{L}^{\prime}_{S4}$ such that the restriction $v_{|\Lambda}$ of $v$ to the domain $\Lambda$ coincides with $\tilde{v}_{p}$.
Proof.
It is an adaptation of the proof of [4, Lemma 3.14]. In this case, more modifications are required because of the conjunction and disjunction connectives, and by the fact of considering a fixed set $\Lambda$.
Suppose first that $\Lambda=For(\Sigma)$. By induction on $n\in\omega$, it will be shown that $\tilde{v}_{p}\in\mathcal{L}^{\prime S4}_{n}$ for every $n$. By Definition 3.15 of partial valuation, $\tilde{v}_{p}\in Val(\mathcal{M}^{\prime}_{S4})={\mathcal{L}^{\prime}}_{0}^{S4}$. Suppose that $\tilde{v}_{p}\in\mathcal{L}^{\prime S4}_{n}$ for a given $n\geq 0$, and let $\alpha$ such that $\mathcal{L}^{\prime S4}_{n}(\alpha)\subseteq\mathcal{D}$. In particular, $\tilde{v}_{p}(\alpha)\in\mathcal{D}$. Suppose that $\tilde{v}_{p}(\alpha)=1$. By Definition 3.18 of partial’ level valuation, there exists a level valuation $w$ in $\mathcal{L}^{\prime}_{S4}$ such that $w(\alpha)=0$. But this is a contradiction, since $w\in\mathcal{L}^{\prime S4}_{n}$. Then $\tilde{v}_{p}(\alpha)=2$ and so $\tilde{v}_{p}\in\mathcal{L}^{\prime S4}_{n+1}$. From this, $\tilde{v}_{p}\in\mathcal{L}^{\prime}_{S4}$ and the result clearly holds.
Suppose now that $\Lambda\neq\emptyset$ is a proper subset of $For(\Sigma)$. Consider an enumeration of $For(\Sigma)$ as in Remark 3.22. The valuation $v$ will be defined by induction on the complexity $n$ of the formulas. Observe that $\Lambda\cup\mathcal{P}\in CS(\Sigma)$ and, for every $n\in\omega$, $\Lambda_{n}\in CS(\Sigma)$ for $\Lambda_{n}:=\Lambda\cup For(\Sigma)_{n}$. In addition, for every $m\in\omega^{2}$ and $n\geq 0$, if $\alpha_{m}$ is such that $\textsf{c}_{1}(\alpha_{m})=n+1$, then $\Lambda_{n}^{m}\in CS(\Sigma)$, where $\Lambda_{n}^{m}=\Lambda_{n}\cup\{\alpha_{m}\}$. Clearly $\Lambda_{n+1}=\cup_{k\in\omega}\Lambda_{n}^{\omega\cdot(n+1)+k}$.
Base $n=0$: Define $v(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$, and $v(p)=0$ if $p$ is a propositional variable which does not belong to $\Lambda$. Clearly, $v\in PLV^{\prime}(\Lambda\cup\mathcal{P})$ up to this stage.
Inductive step: Assume that $v(\alpha)$ was defined for every $\alpha\in\Lambda_{n}$ such that $v(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$, and $v\in PLV^{\prime}(\Lambda_{n})$ up to this stage, for a given $n\geq 0$ (Induction Hypothesis, IH). Let $\alpha=\alpha_{m}$ (for $m\in I_{n+1}$) such that $\alpha\not\in\Lambda$ and $\textsf{c}_{1}(\alpha)=n+1$. Observe that $v(\beta)$ was already defined, for every strict subformula $\beta$ of $\alpha$. Since $v\in PLV^{\prime}(\Lambda_{n})$ then, for every $\delta\in\Lambda_{n}$ such that $v(\delta)=1$ there exists $w\in\mathcal{L}^{\prime}_{S4}$ such that $\mathsf{P}_{\Lambda_{n}}(v,w,\delta)$. Such a $w$ will be denoted by $w^{\delta}$ (observe that it is possible to have more that one $w^{\delta}$ for each $\delta$, and it is possible to have $w^{\delta}=w^{\gamma}$ for $\delta\neq\gamma$). Recall that $\Lambda_{n}^{m}=\Lambda_{n}\cup\{\alpha_{m}\}$ is closed under subformulas. There are four cases to analyze:
[…]
(2.1) There is $\delta\in\Lambda_{n}$ with $v(\delta)=1$ and there exists some $w^{\delta}$ such that $w^{\delta}(\alpha)=0$. Define $v(\alpha)=1$. Hence $v\in PLV^{\prime}(\Lambda_{n}^{m})$ such that $w^{\alpha}=w^{\delta}$.
(2.2) There is $\delta\in\Lambda_{n}$ with $v(\delta)=1$ and there exists some $w^{\delta}$ such that $w^{\delta}(\alpha)=1$. By Lemma 3.20, $v^{\prime}=w^{\delta}_{|\Lambda_{n}^{m}}\in PLV^{\prime}(\Lambda_{n}^{m})$, hence there exists $w^{\prime\prime}\in\mathcal{L}^{\prime}_{S4}$ such that $\mathsf{P}_{\Lambda_{n}^{m}}(v^{\prime},w^{\prime\prime},\alpha)$. Define $v(\alpha)=1$. Hence, $v\in PLV^{\prime}(\Lambda_{n}^{m})$ such that $w^{\alpha}=w^{\prime\prime}$.
(2.3) For every $\delta\in\Lambda_{n}$ such that $v(\delta)=1$, $w^{\delta}(\alpha)=2$ for every $w^{\delta}$. Define $v(\alpha)=2$. Clearly, $v\in PLV^{\prime}(\Lambda_{n}^{m})$.
[…]
(3.1) $\alpha=\beta\to\gamma$. Then $v(\beta)=2$ and $v(\gamma)=1$. Define $v(\alpha)=1$. Then $v\in PLV^{\prime}(\Lambda_{n}^{m})$ such that $w^{\alpha}=w^{\gamma}$. Indeed, since $w^{\gamma}(\beta)=2$ and $w^{\gamma}(\gamma)=0$ then $w^{\gamma}(\alpha)=0$.
(3.2) $\alpha=\beta\land\gamma$. Then $v(\beta)=1$ and $v(\gamma)\in\mathcal{D}$ or vice versa. Then, there exists $w\in\mathcal{L}^{\prime}_{S4}$ (where $w=w^{\beta}$ or $w=w^{\gamma}$) such that $w(\beta)=0$ or $w(\gamma)=0$ and so $w(\alpha)=0$. Define $v(\alpha)=1$. Then $v\in PLV^{\prime}(\Lambda_{n}^{m})$ such that $w^{\alpha}=w^{\beta}$ or $w^{\alpha}=w^{\gamma}$.
[…]
We have shown in cases (1)-(4) how to extend the domain of $v$ to the additional formula $\alpha_{m}$, for every $\alpha_{m}$ such that $\textsf{c}_{1}(\alpha_{m})=n+1$, showing that $v\in PLV^{\prime}(\Lambda_{n}^{m})$. Consider now the simultaneous extension of $v$ with the values assigned to every $\alpha_{m}$ as prescribed above. We obtain so an extension of $v$ to $\Lambda_{n+1}=\cup_{m}\Lambda_{n}^{m}$. By the very definition of partial’ level valuations it follows that $v\in PLV^{\prime}(\Lambda_{n+1})$, and $v(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$. This completes the inductive step. ∎
Proposition 3.24.
Let $\Lambda\in FCS(\Sigma)$. Then $PLV(\Lambda)=PLV^{\prime}(\Lambda)$.
Proof.
Let $\tilde{v}_{p}\in PLV^{\prime}(\Lambda)$. Then, $\tilde{v}_{p}\in PV(\Lambda)$. Suppose that $\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=1$. Then, there exists $w\in\mathcal{L}^{\prime}_{S4}$ such that $\mathsf{P}_{\Lambda}(\tilde{v}_{p},w,\alpha)$. Let $\tilde{w}_{p}=w_{|\Lambda}$. By Lemma 3.20, $\tilde{w}_{p}\in PLV^{\prime}(\Lambda)$ such that $\mathsf{P}_{\Lambda}(\tilde{v}_{p},\tilde{w}_{p},\alpha)$. Now, suppose that there exists $\tilde{w}^{\prime}_{p}\in PLV^{\prime}(\Lambda)$ such that $\mathsf{P}_{\Lambda}(\tilde{v}_{p},\tilde{w}^{\prime}_{p},\alpha)$. By Lemma 3.23, there exists $v^{\prime}\in\mathcal{L}^{\prime}_{S4}$ such that the restriction of $v^{\prime}$ to $\Lambda$ coincides with $\tilde{w}^{\prime}_{p}$. Hence, $\mathsf{P}_{\Lambda}(\tilde{v}_{p},v^{\prime},\alpha)$. This shows that
$\begin{array}[]{lll}PLV^{\prime}(\Lambda)&=&\{\tilde{v}_{p}\in PV(\Lambda)\mid\forall\alpha\in\Lambda\big{(}\tilde{v}_{p}(\alpha)=1\mbox{ implies that }\mathsf{P}_{\Lambda}(\tilde{v}_{p},\tilde{w}_{p},\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $\tilde{w}_{p}\in PLV^{\prime}(\Lambda)$}\big{)}\}.\end{array}$ From this, being $PV(\Lambda)$ finite, it follows that $PLV(\Lambda)=PLV^{\prime}(\Lambda)$. ∎
Theorem 3.25.
Let $\varphi\in For(\Sigma)$, and let $\Lambda$ be the set of subformulas of $\varphi$. Then: $\vdash_{S4}\varphi$ if, and only if, for every partial valuation $\tilde{v}_{p}\in PLV(\Lambda)$ in $\mathcal{M}^{\prime}_{S4}$, $\tilde{v}_{p}(\varphi)=2$.
Proof.
Let $\Lambda$ be the set of subformulas of $\varphi$. Then, $\Lambda\in FCS(\Sigma)$. Suppose that $\vdash_{S4}\varphi$. By Theorem 3.11, $\models_{\mathcal{R}(\mathcal{M}^{\prime}_{S4})}\varphi$. Let $\tilde{v}_{p}\in PLV(\Lambda)$. By Proposition 3.24, $\tilde{v}_{p}\in PLV^{\prime}(\Lambda)$. By Lemma 3.23, there exists $v\in\mathcal{L}^{\prime}_{S4}$ such that the restriction $v_{|\Lambda}$ of $v$ to $\Lambda$ coincides with $\tilde{v}_{p}$. Then, $\tilde{v}_{p}(\varphi)=v(\varphi)=2$, by Corollary 3.11.1.
Conversely, suppose that $\tilde{v}_{p}(\varphi)=2$ for every $\tilde{v}_{p}\in PLV(\Lambda)$. By Proposition 3.24, $\tilde{v}_{p}(\varphi)=2$ for every $\tilde{v}_{p}\in PLV^{\prime}(\Lambda)$. Let $v\in\mathcal{L}^{\prime}_{S4}$. By Lemma 3.20, the restriction $\tilde{v}_{p}:=v_{|\Lambda}$ of $v$ to $\Lambda$ belongs to $PLV^{\prime}(\Lambda)$. Then, $v(\varphi)=\tilde{v}_{p}(\varphi)=2$ and then $\models_{\mathcal{R}(\mathcal{M}^{\prime}_{S4})}\varphi$. By Theorem 3.13, $\vdash_{S4}\varphi$. ∎
This means that Grätz’s algorithm can be adapted to the reduced (and expanded with disjunction and conjunction) RNmatrix $\mathcal{R}(\mathcal{M}^{\prime}_{S4})$.
4 An RNmatrix for IPL
Definition 4.1.
Consider the signature Ω={{¬},{,,}}normal-$\Omega=\{\{\neg\},\{\land,\lor,\rightarrow\}\}$ for intuitionistic propositional logic (IPL). The set of intuitionistic formulas will be denoted by as $For(\Omega)$.
4.1 Gödel-McKinsey-Tarski Box Translation
Definition 4.2.
(Box translation function) Let $f_{\Box}:For(\Omega)\rightarrow For(\Sigma)$ be a function such that:
  1. i)
    $f_{\Box}(p)=\Box p$, if $p$ is a propositional variable;
  2. ii)
    $f_{\Box}(\neg\alpha)=\Box\neg f_{\Box}(\alpha)$;
  3. iii)
[…]
It is widely known that this translation is faithful [9].
4.2 Truth tables for intuitionistic logic
We introduce the truth tables for intuitionistic logic based on the presented translation.
4.2.1 Intuitionistic non-deterministic valuations
Remark 4.3.
If $v$ is a valuation in $\mathcal{M}_{S4}$, then, for every $\alpha\in For(\Sigma)$, $v(\Box\alpha)=\Box^{\mathcal{A}_{S4}}(v(\alpha))$. This follows from the fact that the multioperation $\Box^{\mathcal{A}_{S4}}$ is deterministic.
Definition 4.4.
We define the function $h:For(\Omega)\rightarrow For(\Sigma)$ as follows:
  1. i)
    $h(p)=p$, if $p$ is a propositional variable;
  2. ii)
    $h(\neg\alpha)=\neg f_{\Box}(\alpha)$;
  3. iii)
[…]
We call $h$ as the semi-translation function.
Lemma 4.5.
For every $\alpha\in For(\Omega)$, $f_{\Box}(\alpha)=\Box(h(\alpha))$.
Proof.
Let $\alpha$ be an intuitionistic formula. We proceed by case analysis.
Case $\alpha=p_{n}$
In this case, $f_{\Box}(\alpha)=f_{\Box}(p_{n})=\Box p_{m}$. Also, $\Box(h(\alpha))=\Box(h(p_{n}))=\Box p_{m}$. Note that $n=m$ in both cases. Hence, $f_{\Box}(p_{n})=\Box(h(p_{n}))$.
Case $\alpha=\neg\beta$
By Definition 4.4, $\Box(h(\neg\beta))=\Box(\neg f_{\Box}(\beta))$. Moreover, by Definition 4.2, $f_{\Box}(\neg\beta)=\Box(\neg f_{\Box}(\beta))$. Hence, $f_{\Box}(\neg\beta)=\Box(h(\neg\beta))$.
[…]
By Definition 4.4, we have that for each ##\##, $\Box(h(\beta\#\gamma))=\Box(f_{\Box}(\beta)\#f_{\Box}(\gamma))$. Moreover, by Definition 4.2, $f_{\Box}(\beta\#\gamma)=\Box(f_{\Box}(\beta)\#f_{\Box}(\gamma))$. Therefore, $f_{\Box}(\beta\#\gamma)=\Box(h(\beta\#\gamma))$.
Remark 4.6.
By Lemma 4.5 and Remark 4.3, $v(f_{\Box}(\alpha))=\Box^{\mathcal{A}_{S4}}(v(h(\alpha)))$, for every $\alpha\in For(\Omega)$.
Definition 4.7.
(Truth function) Let $v$ be a valuation function in $\mathcal{M}_{S4}$ and $\alpha\in For(\Omega)$. We define the function $\tilde{v}:For(\Omega)\rightarrow V\times V$ as follows:
$\tilde{v}(\alpha)\triangleq\langle v(h(\alpha));v(f_{\Box}(\alpha))\rangle$
We will say that $\tilde{v}$ is induced by $v$ and will call $\tilde{v}$ as the truth function.
By Remark 4.6, we get $\tilde{v}(\alpha)=\langle v(h(\alpha));\Box^{\mathcal{A}_{S4}}(v(h(\alpha)))\rangle$. This motivates the following:
Definition 4.8 (Intuitionistic truth values).
The set of intuitionistic truth values is given by
$V_{IPL}=\left\{\langle c;\Box^{\mathcal{A}_{S4}}(c)\rangle\mid c\in\mathcal{V}\right\}=\{\mathbf{F},\mathbf{U},\mathbf{T}\}$
where $\mathbf{F}=\langle 0;0\rangle$, $\mathbf{U}=\langle 1;0\rangle$ and $\mathbf{T}=\langle 2;2\rangle$.
Therefore, the truth function 4.7 is a function $\tilde{v}:For(\Omega)\rightarrow\{\mathbf{F},\mathbf{U},\mathbf{T}\}$.
[…]
Definition 4.9 (Intuitionistic multioperations).
Let $\langle a,b\rangle,\langle c,d\rangle\in V_{IPL}$. Define the intuitionistic multioperations as follows:
  1. 1.
    $\neg^{\prime\mathcal{A}_{IPL}}(\langle a;b\rangle)=\left\{\langle c;\Box^{\mathcal{A}_{S4}}(c)\rangle\mid c\in\neg^{\mathcal{A}_{S4}}(b)\right\}$
  2. 2.
    #$\#^{\prime\mathcal{A}_{IPL}}(\langle a;b\rangle,\langle c;d\rangle)=\left\{\langle e;\Box^{\mathcal{A}_{S4}}(e)\rangle\mid e\in\#^{\mathcal{A}^{\prime}_{S4}}(b,d)\right\}$
[…]
Definition 4.10.
The intuitionistic Nmatrix over signature $\Omega$ is given by $\mathcal{M}_{IPL}=\langle V_{IPL},\mathcal{D}_{IPL},\mathcal{O}_{IPL}\rangle$, where $\mathcal{D}_{IPL}=\{\mathbf{T}\}$, $\mathcal{O}_{IPL}=\{\neg^{\mathcal{A}_{IPL}},\rightarrow^{\mathcal{A}_{IPL}},\lor^{\mathcal{A}_{IPL}},\land^{\mathcal{A}_{IPL}}\}$ and each $o\in\mathcal{O}_{IPL}$ is defined in Table 3.
$\neg^{\mathcal{A}_{IPL}}$
[…]
$\rightarrow^{\mathcal{A}_{IPL}}$ F U T
[…]
$\lor^{\mathcal{A}_{IPL}}$ F U T
[…]
$\land^{\mathcal{A}_{IPL}}$ F U T
[…]
Table 3: Tables for #$\#^{\mathcal{A}_{IPL}}$.
Lemma 4.11.
Let #$\#\in\Omega$. Then, #$\#^{\mathcal{A}_{IPL}}=\#^{\prime\mathcal{A}_{IPL}}$. That is, the multioperators in Definition 4.9 coincide with the ones in Definition 4.10.
Proof.
It is straightforward to verify that, for every $z,w\in V_{IPL}$:
  1. 1.
    $\neg^{\mathcal{A}_{IPL}}(z)=\neg^{\prime\mathcal{A}_{IPL}}(z)$; and
  2. 2.
[…]
Definition 4.12.
[…]
Definition 4.13.
Given $n\geq 0$, the set of formulas over $\Omega$ with complexity less or equal than $n$ will be denoted by $For(\Omega)_{n}$. That is: $For(\Omega)_{n}=\{\alpha\in For(\Omega)\mid\textsf{c}(\alpha)\leq n\}$.
Definition 4.14.
Let $Val(\mathcal{M}_{IPL})$ be the set of valuations over $\mathcal{M}_{IPL}$, and let $Val_{+}(\mathcal{M}_{IPL})$ be the set of valuations $v$ in $Val(\mathcal{M}_{IPL})$ such that $v(p)\in\{\mathbf{T},\mathbf{F}\}$ for every propositional variable $p$.
Definition 4.15.
(Level valuations for IPL) Let $Val(\mathcal{M}_{IPL})$ be the set of valuations over $\mathcal{M}_{IPL}$. We define the level $\mathcal{L}^{IPL}_{k}$, where $k\in\mathbb{N}$, as follows:
  1. i)
    $\mathcal{L}_{0}^{IPL}=Val_{+}(\mathcal{M}_{IPL})$;
  2. ii)
    $\mathcal{L}^{IPL}_{k+1}=\{v\in\mathcal{L}^{IPL}_{k}\mid\forall\alpha\in For(\Omega)_{k+1}$, if $v(\alpha)=\mathbf{U}$ then there exists $w\in\mathcal{L}^{IPL}_{k}$ such that $w(\alpha)=\mathbf{F}$ and: $v(\beta)=\mathbf{T}$ implies $w(\beta)=\mathbf{T}$ for every $\beta\in For(\Omega)\}$.
Definition 4.16.
The set of level valuations in $\mathcal{M}_{IPL}$ is defined as the intersection of every level valuation:
$\mathcal{L}_{IPL}=\cap^{\infty}_{k\geq 0}\mathcal{L}^{IPL}_{k}$
Definition 4.17 (RNmatrix for IPL).
Let $\mathcal{M}_{IPL}$ be the Nmatrix defined in Definition 4.10. The RNmatrix for IPL is the pair $\mathcal{R}(\mathcal{M}_{IPL})=\langle\mathcal{M}_{IPL},\mathcal{L}_{IPL}\rangle$.
From now on, $H_{IPL}$ will denote the standard Hilbert calculus for IPL over $\Omega$, given by the following axioms, and with Modus Ponens (MP) as the only inference rule:
$\begin{array}[]{ll}\textsf{(Ax1)}&\alpha\to(\beta\to\alpha)\\[5.69054pt] \textsf{(Ax2)}&(\alpha\to(\beta\to\gamma))\to((\alpha\to\beta)\to(\alpha\to\gamma))\\[5.69054pt] \textsf{(Ax3)}&\alpha\to(\beta\to(\alpha\land\beta))\\[5.69054pt] \textsf{(Ax4)}&(\alpha\land\beta)\to\alpha\\[5.69054pt] \textsf{(Ax5)}&(\alpha\land\beta)\to\beta\\[5.69054pt] \textsf{(Ax6)}&\alpha\to(\alpha\lor\beta)\\[5.69054pt] \textsf{(Ax7)}&\beta\to(\alpha\lor\beta)\\[5.69054pt] \textsf{(Ax8)}&(\alpha\to\gamma)\to((\beta\to\gamma)\to((\alpha\lor\beta)\to\gamma))\\[5.69054pt] \textsf{(Ax9)}&(\beta\to\alpha)\to((\beta\to\neg\alpha)\to\neg\beta)\\[5.69054pt] \textsf{(Ax10)}&\alpha\to(\neg\alpha\to\beta)\\[5.69054pt] \end{array}$
The consequence relation of $H_{IPL}$ will be denoted by $\vdash_{IPL}$. Observe that derivations in $H_{IPL}$ from premises are defined as usual in Hilbert calculi. That is, premises can be used in the derivations in the same way as axioms.
The proof of soundness of $H_{IPL}$ w.r.t. the RNmatrix $\mathcal{R}(\mathcal{M}_{IPL})$ will be postponed to Subsection 4.2.2. We will concentrate now on the completeness proof.
Remark 4.18.
Recall from Subsection 3.2 the notion of $\varphi$-saturated set in L. If $\Delta$ is a $\varphi$-saturated set in the logic generated by $\mathcal{H}_{IPL}$ (that is, IPL) then: $\alpha\in\Delta$ implies that $\neg\alpha\notin\Delta$; $\alpha\vee\beta\in\Delta$ iff $\alpha\in\Delta$ or $\beta\in\Delta$; $\alpha\land\beta\in\Delta$ iff $\alpha\in\Delta$ and $\beta\in\Delta$; $\alpha\to\beta\in\Delta$ and $\alpha\in\Delta$ implies that $\beta\in\Delta$; and $\beta\in\Delta$ implies that $\alpha\to\beta\in\Delta$.
Definition 4.19 (Valuation associated to a $\varphi$-saturated set in IPL).
Let $\Delta\subseteq For(\Omega)$ be a $\varphi$-saturated set in the logic generated by $\mathcal{H}_{IPL}$ (that is, IPL). The valuation associated to $\Delta$ in IPL is the function $v_{\Delta}:For(\Omega)\to V_{IPL}$ defined inductively as follows (here, $p$ is a propositional variable):
[…]
Remark 4.20.
Observe that $v_{\Delta}(\neg\alpha)=\mathbf{U}$ whenever $\alpha$ is undetermined in $\Delta$, that is: $\neg\alpha\notin\Delta$ and $\alpha\notin\Delta$. Hence, the failure of excluded middle (which is a principle valid in classical logic but not in IPL) occurs only for formulas $\alpha$ with such undetermined status in $\Delta$. Analogously, $v_{\Delta}(\alpha\to\beta)=\mathbf{U}$ whenever $\alpha$ and $\alpha\to\beta$ are jointly undetermined in $\Delta$, that is: neither $\alpha$ nor $\alpha\to\beta$ belong to $\Delta$. Hence, the failure of $\alpha\vee(\alpha\to\beta)$ (which is a principle valid in classical logic but not in IPL) occurs only for formulas $\alpha$ and $\alpha\to\beta$ with such joint undetermined status in $\Delta$.
Proposition 4.21.
Let $\Delta\subseteq For(\Omega)$ be a $\varphi$-saturated set in the logic generated by $\mathcal{H}_{IPL}$. Then, the valuation $v_{\Delta}$ associated to $\Delta$ in IPL is a valuation in the Nmatrix $\mathcal{M}_{IPL}$ which belongs to $Val_{+}(\mathcal{M}_{IPL})$ (recall Definition 4.14) such that, for every $\alpha$: $v_{\Delta}(\alpha)=\mathbf{T}$ iff $\alpha\in\Delta$.
Proof.
Observe that, by definition and by Remark 4.18, $v_{\Delta}(\alpha)=\mathbf{T}$ iff $\alpha\in\Delta$, for every formula $\alpha$. Moreover, $v_{\Delta}(p)\in\{\mathbf{T},\mathbf{F}\}$ for every propositional variable $p$.
Negation: Let us prove first that $v_{\Delta}(\neg\alpha)\in\neg^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha))$ for every $\alpha$. If $v_{\Delta}(\neg\alpha)=\mathbf{T}$ then $\neg\alpha\in\Delta$, by definition. Hence, $\alpha\notin\Delta$ and so $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$. From this, $v_{\Delta}(\neg\alpha)=\mathbf{T}\in\{\mathbf{U},\mathbf{T}\}=\neg^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha))$. Now, if $v_{\Delta}(\neg\alpha)=\mathbf{U}$ then $\alpha\notin\Delta$, by definition. Hence $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$ and so $v_{\Delta}(\neg\alpha)=\mathbf{U}\in\{\mathbf{U},\mathbf{T}\}=\neg^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha))$. Finally, if $v_{\Delta}(\neg\alpha)=\mathbf{F}$ then $\alpha\in\Delta$, by definition. Hence $v_{\Delta}(\alpha)=T$ and so $v_{\Delta}(\neg\alpha)=\mathbf{F}\in\{\mathbf{F}\}=\neg^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha))$.
Implication: There are 3 cases:
(I1) Suppose that $v_{\Delta}(\alpha\to\beta)=\mathbf{T}$. Then $\alpha\to\beta\in\Delta$, by definition.
– If $\alpha\notin\Delta$ then $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$. Hence, $v_{\Delta}(\alpha\to\beta)=\mathbf{T}\in{\to}^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
– If $\alpha\in\Delta$ then $\beta\in\Delta$, by MP. Hence $v_{\Delta}(\alpha)=v_{\Delta}(\beta)=\mathbf{T}$ and so $v_{\Delta}(\alpha\to\beta)=\mathbf{T}\in\{\mathbf{T}\}={\to}^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
(I2) Suppose that $v_{\Delta}(\alpha\to\beta)=\mathbf{U}$. Then $\alpha\to\beta\notin\Delta$ and $\alpha\notin\Delta$, by definition. Moreover, $\beta\notin\Delta$, by Remark 4.18. Then $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$ and $v_{\Delta}(\beta)\in\{\mathbf{U},\mathbf{F}\}$. Hence, $v_{\Delta}(\alpha\to\beta)=\mathbf{U}\in\{\mathbf{U},\mathbf{T}\}={\to}^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
(I3) Suppose that $v_{\Delta}(\alpha\to\beta)=\mathbf{F}$. Then $\alpha\to\beta\notin\Delta$ and $\alpha\in\Delta$, by definition. Moreover, $\beta\notin\Delta$, by Remark 4.18. Then $v_{\Delta}(\alpha)=\mathbf{T}$ and $v_{\Delta}(\beta)\in\{\mathbf{U},\mathbf{F}\}$. Hence, $v_{\Delta}(\alpha\to\beta)=\mathbf{F}\in\{\mathbf{F}\}={\to}^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
Conjunction: There are 2 cases:
(C1) Suppose that $v_{\Delta}(\alpha\land\beta)=\mathbf{T}$. Then $\alpha\in\Delta$ and $\beta\in\Delta$, by definition. From this, $v_{\Delta}(\alpha)=v_{\Delta}(\beta)=\mathbf{T}$. Hence, $v_{\Delta}(\alpha\land\beta)=\mathbf{T}\in\{\mathbf{T}\}=\land^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
(C2) Suppose that $v_{\Delta}(\alpha\land\beta)=\mathbf{F}$. Then, either $\alpha\notin\Delta$ or $\beta\notin\Delta$, by definition. From this, $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$ or $v_{\Delta}(\beta)\in\{\mathbf{U},\mathbf{F}\}$. Hence, $v_{\Delta}(\alpha\land\beta)=\mathbf{F}\in\{\mathbf{F}\}=\land^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
Disjunction: There are 2 cases:
(D1) Suppose that $v_{\Delta}(\alpha\vee\beta)=\mathbf{T}$. Then, either $\alpha\in\Delta$ or $\beta\in\Delta$, by definition. From this, $v_{\Delta}(\alpha)=\mathbf{T}$ or $v_{\Delta}(\beta)=\mathbf{T}$. Hence, $v_{\Delta}(\alpha\vee\beta)=\mathbf{T}\in\{\mathbf{T}\}=\vee^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
(D2) Suppose that $v_{\Delta}(\alpha\vee\beta)=\mathbf{F}$. Then $\alpha\notin\Delta$ and $\beta\notin\Delta$, by definition. From this, $v_{\Delta}(\alpha)\in\{\mathbf{U},\mathbf{F}\}$ and $v_{\Delta}(\beta)\in\{\mathbf{U},\mathbf{F}\}$. Hence, $v_{\Delta}(\alpha\land\beta)=\mathbf{F}\in\{\mathbf{F}\}=\vee^{\mathcal{A}_{IPL}}(v_{\Delta}(\alpha),v_{\Delta}(\beta))$.
This concludes the proof. ∎
Proposition 4.22.
Let $\Delta\subseteq For(\Omega)$ be a $\varphi$-saturated set in the logic generated by $\mathcal{H}_{IPL}$. Then, the valuation $v_{\Delta}$ associated to $\Delta$ in IPL is a level valuation in $\mathcal{M}_{IPL}$ such that, for every $\alpha$: $v_{\Delta}(\alpha)=\mathbf{T}$ iff $\alpha\in\Delta$.
Proof.
By Proposition 4.21, $v_{\Delta}(\alpha)=\mathbf{T}$ iff $\alpha\in\Delta$ for every $\alpha$. It will be shown, by induction on $n$, that $v_{\Delta}\in\mathcal{L}^{IPL}_{n}$ for every $n\geq 0$.
Base $n=0$: $v_{\Delta}\in\mathcal{L}^{IPL}_{0}=Val_{+}(\mathcal{M}_{IPL})$, by Proposition 4.21.
[…]
(1) $\alpha=\neg\beta$ for some $\beta$. By Definition 4.19, $\neg\beta\not\in\Delta$ and $\beta\not\in\Delta$. Suppose that $\Delta,\beta\vdash_{IPL}\neg\beta$. By the deduction metatheorem, $\Delta\vdash_{IPL}\beta\to\neg\beta$. By (Ax9) and by MP, $\Delta,\beta\to\beta,\beta\to\neg\beta\vdash_{IPL}\neg\beta$. But $\beta\to\beta\in\Delta$ and $\beta\to\neg\beta\in\Delta$, hence $\Delta\vdash_{IPL}\neg\beta$, that is, $\neg\beta\in\Delta$, a contradiction. From this, $\Delta,\beta\nvdash_{IPL}\neg\beta$. Then, there exists a $\neg\beta$-saturated set $\Delta^{\prime}$ in IPL such that $\Delta\cup\{\beta\}\subseteq\Delta^{\prime}$. By IH, $v_{\Delta^{\prime}}\in\mathcal{L}^{IPL}_{n}$ such that, by Definition 4.19, $v_{\Delta^{\prime}}(\neg\beta)=\mathbf{F}$. Moreover, for every $\delta\in For(\Omega)$, if $v_{\Delta}(\delta)=\mathbf{T}$ then $v_{\Delta^{\prime}}(\delta)=\mathbf{T}$, since $\Delta\subseteq\Delta^{\prime}$.
(2) $\alpha=\beta\to\gamma$ for some $\beta$ and $\gamma$. By Definition 4.19, $\beta\to\gamma\not\in\Delta$ and $\beta\not\in\Delta$. Suppose that $\Delta,\beta\vdash_{IPL}\beta\to\gamma$. By MP, $\Delta,\beta\vdash_{IPL}\gamma$ and so, by the deduction metatheorem, $\Delta\vdash_{IPL}\beta\to\gamma$, a contradiction. From this, $\Delta,\beta\nvdash_{IPL}\beta\to\gamma$. Then, there exists a $(\beta\to\gamma)$-saturated set $\Delta^{\prime\prime}$ in IPL such that $\Delta\cup\{\beta\}\subseteq\Delta^{\prime\prime}$. By IH, $v_{\Delta^{\prime\prime}}\in\mathcal{L}^{IPL}_{n}$ such that, by Definition 4.19, $v_{\Delta^{\prime\prime}}(\beta\to\gamma)=\mathbf{F}$. Moreover, for every $\delta\in For(\Omega)$, if $v_{\Delta}(\delta)=\mathbf{T}$ then $v_{\Delta^{\prime\prime}}(\delta)=\mathbf{T}$, since $\Delta\subseteq\Delta^{\prime\prime}$.
From the previous analysis we conclude that $v_{\Delta}\in\mathcal{L}^{IPL}_{n+1}$. This shows that $v_{\Delta}\in\mathcal{L}^{IPL}_{n}$ for every $n\geq 0$ and so $v_{\Delta}\in\mathcal{L}_{IPL}$. ∎
Theorem 4.23 (Completeness of $\mathcal{H}_{IPL}$ w.r.t. $\mathcal{R}(\mathcal{M}_{IPL})$).
For every $\Gamma$ and $\varphi$: $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$ implies that $\Gamma\vdash_{IPL}\varphi$.
Proof.
Suppose that $\Gamma\nvdash_{IPL}\varphi$. Then, there exists a $\varphi$-saturated set $\Delta$ in IPL such that $\Gamma\subseteq\Delta$. By Proposition 4.22, $v_{\Delta}$ is a level valuation in $\mathcal{M}_{IPL}$ such that $v_{\Delta}(\alpha)=\mathbf{T}$ for every $\alpha\in\Gamma$, but $v_{\Delta}(\varphi)\neq\mathbf{T}$. This shows that $\Gamma\not\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$. ∎
4.2.2 Intuitionistic partial (level) valuations and intuitionistic truth tables
Definition 4.24.
Consider the sets
[…]
Definition 4.25 (Partial valuation in $\mathcal{M}_{IPL}$).
Let $\Lambda\in CS(\Omega)$. A partial valuation in $\mathcal{M}_{IPL}$ is a function $\tilde{v}_{p}:\Lambda\rightarrow V_{IPL}$ such that, for every $\alpha,\beta\in\Lambda$:
  • if $\alpha\in\mathcal{P}\cap\Lambda$ then $\tilde{v}_{p}(\alpha)\in\{\mathbf{T},\mathbf{F}\}$;
  • if $\neg\alpha\in\Lambda$ then $\tilde{v}_{p}(\neg\alpha)\in\neg^{\mathcal{A}_{IPL}}(\tilde{v}_{p}(\alpha))$;
[…]
Let $iPV(\Lambda)$ be the set of partial valuations in $\mathcal{M}_{IPL}$ with domain $\Lambda$.
Definition 4.26 (Intuitionistic partial level valuation over $\Lambda$).
Let $\Lambda\in FCS(\Omega)$. A partial valuation $\tilde{v}_{p}\in iPV(\Lambda)$ is an intuitionistic partial level valuation in $\mathcal{M}_{IPL}$ over $\Lambda$ iff
  1. $\forall\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=\mathbf{U}$, there exists an intuitionistic partial level valuation $\tilde{w}_{p}$ in $\mathcal{M}_{IPL}$ over $\Lambda$ such that $\tilde{w}_{p}(\alpha)=\mathbf{F}$ and, $\forall\beta\in\Lambda$, $\tilde{w}_{p}(\beta)=\mathbf{T}$ whenever $\tilde{v}_{p}(\beta)=bT$.
The set of intuitionistic partial level valuations in $\mathcal{M}_{IPL}$ over $\Lambda$ will be denoted by $iPLV(\Lambda)$.
Remark 4.27.
Given $v,w$ and $\alpha$ let $\mathsf{P}_{\Lambda}^{i}(v,w,\alpha)$ iff $w(\alpha)=\mathbf{F}$ and, $\forall\beta\in\Lambda$, $w(\beta)=\mathbf{T}$ whenever $v(\beta)=\mathbf{T}$. If $\Lambda\in FCS(\Omega)$ then
$\begin{array}[]{lll}iPLV(\Lambda)&=&\{\tilde{v}_{p}\in iPV(\Lambda)\mid\forall\alpha\in\Lambda\big{(}\tilde{v}_{p}(\alpha)=\mathbf{U}\mbox{ implies that }\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},\tilde{w}_{p},\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $\tilde{w}_{p}\in iPLV(\Lambda)$}\big{)}\}.\end{array}$
As observed in Remark 3.7 for $\mathcal{M}_{S4}$, since $\Lambda$ is finite then $iPV(\Lambda)$ is finite, hence the definition of $iPLV(\Lambda)$ is not cyclic.
Definition 4.28.
An intuitionistic truth table in $\mathcal{M}_{IPL}$ is an exhaustive list of all intuitionistic partial level valuations for a given finite domain closed under subformulas.
Remark 4.29.
Each of these partial valuations corresponds to a row of the table.
The following is an example of the algorithm for searching the (intuitionistic) partial level valuations. [^5] This example, as the ones presented in Appendix A, were automatically generated by a Coq module for RNmatrices written by R. Leme, see  https://github.com/renatoleme/Forest We start with the full table for $\neg\neg(p\lor\neg p)$.
Table 4: Initial table.
Row (ID) $p$$\neg p$ $p\lor\neg p$$\neg(p\lor\neg p)$$\neg\neg(p\lor\neg p)$
(1) F U F U U
(2) F U F U T
(3) F U F T F
(4) F T T F U
(5) F T T F T
(6) T F T F U
(7) T F T F T
Table 5: First cycle.
Row (ID) $p$$\neg p$ $p\lor\neg p$$\neg(p\lor\neg p)$$\neg\neg(p\lor\neg p)$
(1) $\times$ 6, 7 $\times$ 4, 5, 6, 7 3
(2) $\times$ 7 $\times$ 5, 7 $\times$
(3) $\times$ $\emptyset$ $\times$ $\times$ $\times$
(4) $\times$ $\times$ $\times$ $\times$ $\emptyset$
(5) $\times$ $\times$ $\times$ $\times$ $\times$
(6) $\times$ $\times$ $\times$ $\times$ $\emptyset$
(7) $\times$ $\times$ $\times$ $\times$ $\times$
The cells marked with the multiplication symbol ($\times$) in a collumn $\alpha$ of a partial valuation $v$ indicates that $v(\alpha)\neq\mathbf{U}$, so we can ignore that cell. The cells marked with the empty set symbol ($\emptyset$) in a collumn $\alpha$ of a partial valuation $v$ indicates that $v(\alpha)=\mathbf{U}$ and there is no other line $w$ to validate $v$, so that line must be removed.
Table 6: Intermediate table.
Row (ID) $p$$\neg p$ $p\lor\neg p$$\neg(p\lor\neg p)$$\neg\neg(p\lor\neg p)$
(1) F U F U U
(2) F U F U T
(5) F T T F T
(7) T F T F T
Table 7: Second cycle.
Row (ID) $p$$\neg p$ $p\lor\neg p$$\neg(p\lor\neg p)$$\neg\neg(p\lor\neg p)$
(1) $\times$ 7 $\times$ 5, 7 $\emptyset$
(2) $\times$ 7 $\times$ 5, 7 $\times$
(5) $\times$ $\times$ $\times$ $\times$ $\times$
(7) $\times$ $\times$ $\times$ $\times$ $\times$
Table 8: Final table.
Row (ID) $p$$\neg p$ $p\lor\neg p$$\neg(p\lor\neg p)$$\neg\neg(p\lor\neg p)$
(2) F U F U T
(5) F T T F T
(7) T F T F T
Table 9: Validators.
Row (ID) $p$$\neg p$ $p\lor\neg p$$\neg(p\lor\neg p)$$\neg\neg(p\lor\neg p)$
(2) $\times$ 7 $\times$ 5, 7 $\times$
(5) $\times$ $\times$ $\times$ $\times$ $\times$
(7) $\times$ $\times$ $\times$ $\times$ $\times$
Note that, in the validators table, there is no occurrence of $\emptyset$. So, there is no further row to remove and we can stop the searching algorithm.
4.2.3 Soundness and completeness of the intuitionistic truth tables
Now, it will be shown that the generalized notion of truth tables for IPL proposed in Subsection 4.2.2 is sound and complete. As in the case of S4, this will be shown by proving first analyticity and co-analyticity w.r.t. level valuations in $\mathcal{M}_{IPL}$. In order to do this it is considered, adapting [4, Definition 3.12], the intermediary notion of intuitionistic partial’ level valuation in $\mathcal{R}(\mathcal{M}_{IPL})$ over $\Lambda$.
Definition 4.30 (Intuitionistic partial’ level valuation over $\Lambda$).
Let $\Lambda\in CS(\Omega)$. An intuitionistic partial valuation $\tilde{v}_{p}\in iPV(\Lambda)$ is an intuitionistic partial’ level valuation in $\mathcal{R}(\mathcal{M}_{IPL})$ over $\Lambda$ iff
  1. $\forall\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=\mathbf{U}$, there exists a level valuation $w$ in $\mathcal{L}_{IPL}$ such that $w(\alpha)=\mathbf{F}$ and, $\forall\beta\in\Lambda$, $w(\beta)=\mathbf{T}$ whenever $\tilde{v}_{p}(\beta)=\mathbf{T}$.
The set of intuitionistic partial’ level valuations in $\mathcal{R}(\mathcal{M}_{IPL})$ over $\Lambda$ will be denoted by $iPLV^{\prime}(\Lambda)$.
Remark 4.31.
Recall the predicate $\mathsf{P}_{\Lambda}^{i}(v,w,\alpha)$ introduced in Remark 4.27. Clearly,
$\begin{array}[]{lll}iPLV^{\prime}(\Lambda)&=&\{\tilde{v}_{p}\in iPV(\Lambda)\mid\forall\alpha\in\Lambda\big{(}\tilde{v}_{p}(\alpha)=\mathbf{U}\mbox{ implies that }\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},w,\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $w\in\mathcal{L}_{IPL}$}\big{)}\}.\end{array}$
Lemma 4.32 (Co-analyticity lemma for $\mathcal{R}(\mathcal{M}_{IPL})$).
Let $v\in\mathcal{L}_{IPL}$ and $\Lambda\in CS(\Omega)$. Then, the restriction $\tilde{v}_{p}:=v_{|\Lambda}$ of $v$ to the domain $\Lambda$ belongs to $iPLV^{\prime}(\Lambda)$.
Proof.
It is an adaptation of the proof of [4, Lemma 3.13]. Suppose, by contradiction, that $\tilde{v}_{p}\notin iPLV^{\prime}(\Lambda)$. Let $\Gamma:=\{\beta\in\Lambda\mid\tilde{v}_{p}(\beta)=\mathbf{T}\}$. Then, there exists $\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=\mathbf{U}$ and, for every $w\in\mathcal{L}_{IPL}$, $w(\alpha)=\mathbf{F}$ implies that $w(\beta)\neq\mathbf{T}$ for some $\beta\in\Gamma$. By contraposition, this means that $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha$. But $v(\beta)=\mathbf{T}$ for every $\beta\in\Gamma$ and $v(\alpha)=\mathbf{U}$, a contradiction. From this we infer that $\tilde{v}_{p}\in iPLV^{\prime}(\Lambda)$. ∎
Remark 4.33.
Observe that, as it was done in Remark 3.22 for S4, it is possible to define an enumeration $\alpha_{1},\alpha_{2},\ldots\alpha_{m}\ldots$ (for $m\in\omega^{2}$) of $For(\Omega)$ such that $\textsf{c}(\alpha_{i})\leq\textsf{c}(\alpha_{j})$ if $i\leq j$ and, for every $i$ such that $\textsf{c}(\alpha_{i})>0$, if $\beta$ is a strict subformula of $\alpha_{i}$ then $\beta=\alpha_{j}$ for some $j<i$. Such an enumeration of $For(\Omega)$ can be defined as follows: every formula have and index in $\omega^{2}$, which is a denumerable ordinal, such that all the formulas with complexity 00 (which form a denumerable set) are placed first, with indexes in $I_{0}:=\omega$; after this, all the formulas with complexity $1$ (which form a denumerable set) are placed with an index in $I_{1}:=\omega\cdot 2\setminus\omega=\{\omega,\omega+1,\ldots\}$; in general, the formulas with complexity $n$ (which form a denumerable set) have an index in $I_{n}:=\omega\cdot(n+1)\setminus\omega\cdot n=\{\omega\cdot n,\omega\cdot n+1,\ldots\}$.
Lemma 4.34 (Analyticity lemma for $\mathcal{R}(\mathcal{M}_{IPL})$).
Let $\Lambda\in CS(\Omega)$ and $\tilde{v}_{p}\in iPLV^{\prime}(\Lambda)$. Then, there exists $v\in\mathcal{L}_{IPL}$ such that the restriction $v_{|\Lambda}$ of $v$ to the domain $\Lambda$ coincides with $\tilde{v}_{p}$.
Proof.
It is an adaptation of the proof of [4, Lemma 3.14] to $\mathcal{R}(\mathcal{M}_{IPL})$.
Suppose first that $\Lambda=For(\Omega)$. By induction on $n\in\omega$, it will be shown that $\tilde{v}_{p}\in\mathcal{L}^{IPL}_{n}$ for every $n$. By Definition 4.25 of partial valuation in $\mathcal{M}_{IPL}$, $\tilde{v}_{p}\in Val_{+}(\mathcal{M}_{IPL})={\mathcal{L}}_{0}^{IPL}$. Suppose that $\tilde{v}_{p}\in\mathcal{L}^{IPL}_{n}$ for a given $n\geq 0$, and let $\alpha\in For(\Omega)_{n+1}$ such that $\tilde{v}_{p}(\alpha)=\mathbf{U}$. By Definition 4.30 of intuitionistic partial’ level valuation, there exists a level valuation $w$ in $\mathcal{L}_{IPL}$ (and so, $w\in\mathcal{L}^{IPL}_{n}$) such that $\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},w,\alpha)$. Since $\Lambda=For(\Omega)$, this means that $\tilde{v}_{p}\in\mathcal{L}^{IPL}_{n+1}$. From this, $\tilde{v}_{p}\in\mathcal{L}_{IPL}$ and the result clearly holds.
Suppose now that $\Lambda\neq\emptyset$ is a proper subset of $For(\Omega)$. Consider an enumeration of $For(\Omega)$ as in Remark 4.33. The valuation $v$ will be defined by induction on the complexity $n$ of the formulas. Observe that $\Lambda\cup\mathcal{P}\in CS(\Omega)$ and, for every $n\in\omega$, $\bar{\Lambda}_{n}\in CS(\Omega)$ for $\bar{\Lambda}_{n}:=\Lambda\cup For(\Omega)_{n}$. In addition, for every $m\in\omega^{2}$ and $n\geq 0$, if $\alpha_{m}$ is such that $\textsf{c}(\alpha_{m})=n+1$, then $\bar{\Lambda}_{n}^{m}\in CS(\Sigma)$, where $\bar{\Lambda}_{n}^{m}=\bar{\Lambda}_{n}\cup\{\alpha_{m}\}$. Clearly $\bar{\Lambda}_{n+1}=\cup_{k\in\omega}\bar{\Lambda}_{n}^{\omega\cdot(n+1)+k}$.
Base $n=0$: Define $v(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$, and $v(p)=\mathbf{F}$ if $p$ is a propositional variable which does not belong to $\Lambda$. Clearly, $v\in PLV^{\prime}(\Lambda\cup\mathcal{P})$ up to this stage.
Inductive step: Assume that $v(\alpha)$ was defined for every $\alpha\in\bar{\Lambda}_{n}$ such that $v(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$, and $v\in iPLV^{\prime}(\bar{\Lambda}_{n})$ up to this stage, for a given $n\geq 0$ (Induction Hypothesis, IH). Let $\alpha=\alpha_{m}$ (for $m\in I_{n+1}$) such that $\alpha\not\in\Lambda$ and $\textsf{c}(\alpha)=n+1$. Observe that $v(\beta)$ was already defined, for every strict subformula $\beta$ of $\alpha$. Since $v\in iPLV^{\prime}(\bar{\Lambda}_{n})$ then, for every $\delta\in\bar{\Lambda}_{n}$ such that $v(\delta)=\mathbf{U}$ there exists $w\in\mathcal{L}_{IPL}$ such that $\mathsf{P}_{\bar{\Lambda}_{n}}^{i}(v,w,\delta)$. Such a $w$ will be denoted by $w^{\delta}$ (observe that it is possible to have more that one $w^{\delta}$ for each $\delta$, and it is possible to have $w^{\delta}=w^{\gamma}$ for $\delta\neq\gamma$). Recall that $\bar{\Lambda}_{n}^{m}=\bar{\Lambda}_{n}\cup\{\alpha_{m}\}$ is closed under subformulas. There are three cases to analyze:
(1) $\alpha=\neg\beta$ such that $\neg^{\mathcal{A}_{IPL}}v(\beta)=\{\mathbf{U},\mathbf{T}\}$, or $\alpha=\beta\to\gamma$ such that $\to^{\mathcal{A}_{IPL}}(v(\beta),v(\gamma))=\{\mathbf{U},\mathbf{T}\}$. There are three subcases to analyze:
(1.1) There is $\delta\in\bar{\Lambda}_{n}$ with $v(\delta)=\mathbf{U}$ and there exists some $w^{\delta}$ such that $w^{\delta}(\alpha)=\mathbf{F}$. Define $v(\alpha)=\mathbf{U}$. Hence $v\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m})$ such that $w^{\alpha}=w^{\delta}$.
(1.2) There is $\delta\in\bar{\Lambda}_{n}$ with $v(\delta)=\mathbf{U}$ and there exists some $w^{\delta}$ such that $w^{\delta}(\alpha)=\mathbf{U}$. By Lemma 4.32, $v^{\prime}=w^{\delta}_{|\bar{\Lambda}_{n}^{m}}\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m})$, hence there exists $w^{\prime\prime}\in\mathcal{L}_{IPL}$ such that $\mathsf{P}_{\bar{\Lambda}_{n}^{m}}^{i}(v^{\prime},w^{\prime\prime},\alpha)$. Define $v(\alpha)=\mathbf{U}$. Hence, $v\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m})$ such that $w^{\alpha}=w^{\prime\prime}$.
(1.3) For every $\delta\in\bar{\Lambda}_{n}$ such that $v(\delta)=\mathbf{U}$, $w^{\delta}(\alpha)=\mathbf{T}$ for every $w^{\delta}$. Define $v(\alpha)=\mathbf{T}$. Clearly, $v\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m})$.
[…]
We have shown in cases (1)-(3) how to extend the domain of $v$ to the additional formula $\alpha_{m}$, for every $\alpha_{m}$ such that $\textsf{c}(\alpha_{m})=n+1$, showing that $v\in iPLV^{\prime}(\bar{\Lambda}_{n}^{m})$. Consider now the simultaneous extension of $v$ with the values assigned to every $\alpha_{m}$ as prescribed above. We obtain so an extension of $v$ to $\bar{\Lambda}_{n+1}=\cup_{m}\bar{\Lambda}_{n}^{m}$. By the very definition of intuitionistic partial’ level valuations it follows that $v\in iPLV^{\prime}(\bar{\Lambda}_{n+1})$, and $v(\alpha)=\tilde{v}_{p}(\alpha)$ if $\alpha\in\Lambda$. This completes the inductive step. ∎
Proposition 4.35.
Let $\Lambda\in FCS(\Omega)$. Then $iPLV(\Lambda)=iPLV^{\prime}(\Lambda)$.
Proof.
Let $\tilde{v}_{p}\in iPLV^{\prime}(\Lambda)$. Then, $\tilde{v}_{p}\in iPV(\Lambda)$. Suppose that $\alpha\in\Lambda$ such that $\tilde{v}_{p}(\alpha)=\mathbf{U}$. Then, there exists $w\in\mathcal{L}_{IPL}$ such that $\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},w,\alpha)$. Let $\tilde{w}_{p}=w_{|\Lambda}$. By Lemma 4.32, $\tilde{w}_{p}\in iPLV^{\prime}(\Lambda)$ such that $\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},\tilde{w}_{p},\alpha)$. Now, suppose that there exists $\tilde{w}^{\prime}_{p}\in iPLV^{\prime}(\Lambda)$ such that $\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},\tilde{w}^{\prime}_{p},\alpha)$. By Lemma 4.34, there exists $v^{\prime}\in\mathcal{L}_{IPL}$ such that the restriction of $v^{\prime}$ to $\Lambda$ coincides with $\tilde{w}^{\prime}_{p}$. Hence, $\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},v^{\prime},\alpha)$. This shows that
$\begin{array}[]{lll}iPLV^{\prime}(\Lambda)&=&\{\tilde{v}_{p}\in PV(\Lambda)\mid\forall\alpha\in\Lambda\big{(}\tilde{v}_{p}(\alpha)=\mathbf{U}\mbox{ implies that }\mathsf{P}_{\Lambda}^{i}(\tilde{v}_{p},\tilde{w}_{p},\alpha)\\[2.84526pt] &&\hskip 19.91692pt\mbox{ for some $\tilde{w}_{p}\in iPLV^{\prime}(\Lambda)$}\big{)}\}.\end{array}$ From this, being $iPV(\Lambda)$ finite, it follows that $iPLV(\Lambda)=iPLV^{\prime}(\Lambda)$. ∎
Definition 4.36.
Let $\Gamma\cup\{\varphi\}$ be a finite subset of $For(\Omega)$. Let $\Lambda$ be the (finite) set of subformulas of $\Gamma\cup\{\varphi\}$. We say that $\varphi$ is a consequence of $\Gamma$ w.r.t. intuitionistic truth-tables, denoted by $\Gamma\models_{iPLV}\varphi$, if, for every intuitionistic partial level valuation $\tilde{v}_{p}\in iPLV(\Lambda)$, if $\tilde{v}_{p}(\beta)=\mathbf{T}$ for every $\beta\in\Gamma$ then $\tilde{v}_{p}(\varphi)=\mathbf{T}$. In particular, $\emptyset\models_{iPLV}\varphi$, denoted simply by $\models_{iPLV}\varphi$, if $\tilde{v}_{p}(\varphi)=\mathbf{T}$ for every $\tilde{v}_{p}\in iPLV(\Lambda)$, where $\Lambda$ is the set of subformulas of $\varphi$.
Theorem 4.37 (Soundness and completeness of intuitionistic truth-tables w.r.t. $\mathcal{R}(\mathcal{M}_{IPL})$).
Let $\Gamma\cup\{\varphi\}$ be a finite subset of $For(\Omega)$. Then: $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$ if, and only if, $\Gamma\models_{iPLV}\varphi$.
Proof.
Let $\Lambda$ be the set of subformulas of $\Gamma\cup\{\varphi\}$. Observe that $\Lambda\in FCS(\Omega)$. Suppose that $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$. Let $\tilde{v}_{p}\in iPLV(\Lambda)$ such that $\tilde{v}_{p}(\beta)=\mathbf{T}$ for every $\beta\in\Gamma$. By Proposition 4.35, $\tilde{v}_{p}\in iPLV^{\prime}(\Lambda)$. By Lemma 4.34, there exists $v\in\mathcal{L}_{IPL}$ such that the restriction $v_{|\Lambda}$ of $v$ to $\Lambda$ coincides with $\tilde{v}_{p}$. Then, $v(\beta)=\mathbf{T}$ for every $\beta\in\Gamma$ and so, by hypothesis, $v(\varphi)=\mathbf{T}$. That is, $\tilde{v}_{p}(\varphi)=\mathbf{T}$. This shows that $\Gamma\models_{iPLV}\varphi$.
Conversely, suppose that $\Gamma\models_{iPLV}\varphi$. Let $v\in\mathcal{L}_{IPL}$ such that $v(\beta)=\mathbf{T}$ for every $\beta\in\Gamma$. By Lemma 4.32, the restriction $\tilde{v}_{p}:=v_{|\Lambda}$ of $v$ to $\Lambda$ belongs to $iPLV^{\prime}(\Lambda)$. By Proposition 4.35, $\tilde{v}_{p}\in iPLV(\Lambda)$ such that $\tilde{v}_{p}(\beta)=\mathbf{T}$ for every $\beta\in\Gamma$. By hypothesis, $\tilde{v}_{p}(\varphi)=\mathbf{T}$. Then, $v(\varphi)=\mathbf{T}$ and so $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$. ∎
Proposition 4.38.
Let $\Gamma\cup\{\alpha,\beta\}\subseteq For(\Omega)$. If $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha\to\beta$ then $\Gamma,\alpha\models_{\mathcal{R}(\mathcal{M}_{IPL})}\beta$.
Proof.
Assume that $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha\to\beta$, and let $v\in\mathcal{L}_{IPL}$ such that $v(\gamma)=v(\alpha)=\mathbf{T}$ for every $\gamma\in\Gamma$. By hypothesis, $v(\alpha\to\beta)=\mathbf{T}$. By definition of $\to^{\mathcal{A}_{IPL}}$, and given that $v\in Val(\mathcal{M}_{IPL})$, it follows from this that $v(\beta)=\mathbf{T}$. This shows that $\Gamma,\alpha\models_{\mathcal{R}(\mathcal{M}_{IPL})}\beta$. ∎
Theorem 4.39 (Soundness of $\mathcal{H}_{IPL}$ w.r.t. $\mathcal{R}(\mathcal{M}_{IPL})$).
For every $\Gamma$ and $\varphi$: $\Gamma\vdash_{IPL}\varphi$ implies that $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$.
Proof.
Since the consequence relation $\mathcal{H}_{IPL}$ of the Hilbert calculus for IPL is finitary and satisfies the deduction metatheorem, and taking into account Proposition 4.38, it is enough to prove the result for $\Gamma=\emptyset$. We prove first the following two facts:
Fact 1: If $\varphi$ is an instance of an axiom in $\mathcal{H}_{IPL}$ then $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$. Indeed, let $\varphi$ be an instance of an axiom in $\mathcal{H}_{IPL}$, and let $\Lambda$ be the set of subformulas of $\varphi$. Then, $\Lambda\in FCS(\Omega)$. It is routine to check that the truth-table for $\varphi$ over $\Lambda$ validates $\varphi$ (see Appendix A). That is: for every intuitionistic partial level valuation $\tilde{v}_{p}\in iPLV(\Lambda)$, $\tilde{v}_{p}(\varphi)=\mathbf{T}$. By Theorem 4.37 it follows that $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$.
Fact 2: If $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha$ and $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha\to\beta$ then $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\beta$. Indeed, assume that $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha$ and $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha\to\beta$. If $v\in\mathcal{L}_{IPL}$ then $v(\alpha)=v(\alpha\to\beta)=\mathbf{T}$. By definition of $\to^{\mathcal{A}_{IPL}}$, and by the fact that $v\in Val(\mathcal{M}_{IPL})$, it follows that $v(\beta)=\mathbf{T}$. This means that $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\beta$.
Now, suppose that $\vdash_{IPL}\varphi$. By induction on the length of a derivation of $\varphi$ in $\mathcal{H}_{IPL}$, and by Fact 1 and Fact 2 it is immediate to prove that $\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$. ∎
Corollary 4.39.1 (Finitariness for $\mathcal{R}(\mathcal{M}_{IPL})$).
Let $\Gamma\cup\{\varphi\}\subseteq For(\Omega)$. If $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$ then $\Gamma_{0}\models_{\mathcal{R}(\mathcal{M}_{IPL})}\varphi$ for some finite subset $\Gamma_{0}$ of $\Gamma$.
Proof.
It follows from Theorems 4.23, 4.39 and the fact that $\mathcal{H}_{IPL}$ is finitary. ∎
Corollary 4.39.2 (Deduction metatheorem for $\mathcal{R}(\mathcal{M}_{IPL})$).
Let $\Gamma\cup\{\alpha,\beta\}\subseteq For(\Omega)$. If $\Gamma,\alpha\models_{\mathcal{R}(\mathcal{M}_{IPL})}\beta$ then $\Gamma\models_{\mathcal{R}(\mathcal{M}_{IPL})}\alpha\to\beta$.
Proof.
It follows from Theorems 4.23, 4.39 and the fact that $\mathcal{H}_{IPL}$ satisfies the deduction metatheorem. ∎
Corollary 4.39.3 (Deduction-detachment metatheorem for intuitionistic truth-tables).
Let $\Gamma\cup\{\alpha,\beta\}$ be a finite subset of $For(\Omega)$. Then, $\Gamma,\alpha\models_{iPLV}\beta$ if, and only if, $\Gamma\models_{iPLV}\alpha\to\beta$.
Proof.
It follows from Theorem 4.37, Proposition 4.38 and Corollary 4.39.2. ∎
It is worth noting that the latter result is useful to reduce the complexity of the truth-tables: instead of generating the (reduced) truth table for a formula $\delta=\beta_{1}\to(\beta_{2}\to(\ldots\to(\beta_{n}\to\varphi)\ldots))$ to be tested as a tautology, it suffices generating the (reduced) truth-table for $\{\beta_{1},\ldots,\beta_{n},\varphi\}$ and then determine if, in every row such that $\beta_{1},\ldots,\beta_{n}$ receive simultaneously the value $\mathbf{T}$, $\varphi$ also receives the value $\mathbf{T}$. This reduces the number of rows required to test $\delta$.
Corollary 4.39.4 (Soundness and completeness of $\mathcal{H}_{IPL}$ w.r.t. intuitionistic truth-tables).
Let $\Gamma\cup\{\varphi\}$ be a finite subset of $For(\Omega)$. Then: $\Gamma\vdash_{IPL}\varphi$ if, and only if, $\Gamma\models_{iPLV}\varphi$.
Proof.
It follows from Theorems 4.39, 4.23 and 4.37. ∎
The table is never empty.
Lemma 4.40.
Let $\Delta\in CS(\Sigma)$ and $\beta\in For(\Sigma)$. If $v_{p}:\Delta\to\mathcal{V}$ is $iPLV$ and $sub(\beta)\subseteq\Delta$, then $v_{p}$ has at least one extension $v^{\prime}_{p}:\Delta\cup\{\beta\}\to\mathcal{V}$ such that $v^{\prime}_{p}$ is $iPLV$.
Proof.
By hypothesis, $v_{p}$ is iPLV. Then, by analiticity of $\mathcal{R}(\mathcal{M}_{IPL})$, there is a level valuation $v:For(\Sigma)\to\mathcal{V}$ such that $v_{p}=v_{\mid\Delta}$, i.e, for every $\alpha\in\Delta$, $v(\alpha)=v_{p}(\alpha)$. More then that, $v$ is defined for $\beta$. Now, given that $sub(\beta)\subseteq\Delta$ and $\Delta\in CS(\Sigma)$, then $\Delta\cup\{\beta\}\in CS(\Sigma)$ and, by co-analiticity of $\mathcal{R}(\mathcal{M}_{IPL})$, $v^{\prime}_{p}=v_{\mid\Delta\cup\{\beta\}}$ is iPLV. ∎
5 Complexity
Given a non-deterministic matrix $\mathcal{M}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$ defined over a language $\mathcal{L}$ and a formula $\alpha\in\mathcal{L}$, we can estimate the total number of rows of the final truth table for $\alpha$ via the following upper bound.
Let $sub(\alpha)\subseteq\mathcal{L}$ be the set of all sub-formulas of $\alpha$, $atoms(\alpha)\subseteq sub(\alpha)$ the set of sub-formulas whose complexity is 00, and $\kappa\geq 1$ the branching factor. Then, the maximum number of rows of the truth table for $\alpha$ in $\mathcal{M}$ is given by
Definition 5.1.
$ub_{\mathcal{M}}(\alpha)=\kappa^{(|sub(\alpha)|-|atoms(\alpha)|)}\times|\mathcal{V}|^{|atoms(\alpha)|}$ (1)
In 5.1, we assume that the branching factor is the maximum degree of non-determinism. More formally,
Definition 5.2.
Let $\mathcal{M}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$ be some nmatrix. The branching factor of $\mathcal{M}$ is the highest $\kappa$ such that, for every n-ary ~$\tilde{\circ}\in\mathcal{O}$, every formula $\alpha_{1},\ldots,\alpha_{n}$ and every valuation $v\in Val(\mathcal{M})$, |~$|\tilde{\circ}(v(\alpha_{1}),\ldots,v(\alpha_{n}))|\leq\kappa$.
For example, in the case of $\mathcal{M}_{IPL}$, we have that disjunction and conjunction have branching factor $1$, while implication and negation have branching factor $2$. Therefore, in this case, $\kappa=2$.
Remark 5.3.
If we consider the valuations in the nmatrix as branches in trees, we can say that $|\mathcal{V}|^{|atoms(\alpha)|}$ gives the number of trees, while $\kappa^{(|sub(\alpha)|-|atoms(\alpha)|)}$ gives the (maximum) number of branches of each tree.
We can also obtain a lower bound by assuming $\kappa=1$, which is the branching factor of deterministic matrices. This gives us
Definition 5.4.
$lb_{\mathcal{M}}(\alpha)=|\mathcal{V}|^{|atoms(\alpha)|}$ (2)
From (5.1) and (5.4), it follows that
$ub_{\mathcal{M}}(\alpha)=\kappa^{(|sub(\alpha)|-|atoms(\alpha)|)}\times lb_{\mathcal{M}}(\alpha)$ (3)
$p$ $p$
$\mathcal{V}_{1}$ $\mathcal{V}_{1}$
$\vdots$ $\vdots$
$\mathcal{V}_{n}$ $\mathcal{V}_{n}$
Table 10: Base case.
Now we shall prove that these are faithful lower and upper bounds for the number of rows.
Theorem 5.5.
Let $\mathcal{M}$ be a nmatrix and $rows(\alpha)$ the set of every row of the truth table for some formula $\alpha$. Then
$lb_{\mathcal{M}}(\alpha)\leq|rows(\alpha)|\leq ub_{\mathcal{M}}(\alpha)$
Proof.
Let $\mathcal{M}=\langle\mathcal{V},\mathcal{D},\mathcal{O}\rangle$ be the nmatrix for the language of $\alpha$, where $\mathcal{V}=\{\mathcal{V}_{1},\ldots,\mathcal{V}_{n}\}$. Induction on the complexity of $\alpha$.
Base case If $\alpha=p$, the truth table is depicted in Table 10, which has $|\mathcal{V}|=n$ rows. Now, $|sub(\alpha)|=|atoms(\alpha)|=1$. For this reason, $lb_{\mathcal{M}}(\alpha)=|\mathcal{V}|^{1}=|\mathcal{V}|$ and $ub(\alpha)=\kappa^{0}\times|\mathcal{V}|^{1}=|\mathcal{V}|$. Therefore, $lb_{\mathcal{M}}(p)=|rows(p)|=ub_{\mathcal{M}}(p)$.
IH If $c(\alpha)<c(\alpha^{\prime})$, then the theorem holds for $\alpha$.
$\alpha=\neg\beta$
Let $rows(\beta)=\{v_{0},\ldots,v_{n}\}$ be the list of rows of the nmatrix for $\beta$. Because $atoms(\beta)=atoms(\neg\beta)$, we conclude that $lb_{\mathcal{M}}(\neg\beta)=lb_{\mathcal{M}}(\beta)$. Now, observe that $|sub(\neg\beta)|=|sub(\beta)|+1$. From this, we get $ub_{\mathcal{M}}(\neg\beta)=\kappa^{|sub(\beta)|-|atoms(\neg\beta)|+1}\times lb_{\mathcal{M}}(\neg\beta)=ub_{\mathcal{M}}(\beta)\times\kappa$. By the definition of branching factor, for every $v_{n}\in rows(\beta)$, |¬~$|\tilde{\neg}(v_{n}(\beta))|\leq\kappa$. Therefore, $|rows(\neg\beta)|\leq|rows(\beta)|\times\kappa$. Now, by (IH), $lb_{\mathcal{M}}(\beta)\leq|rows(\beta)|\leq ub_{\mathcal{M}}(\beta)$. Therefore, $lb_{\mathcal{M}}(\neg\beta)\leq|rows(\neg\beta)|\leq ub_{\mathcal{M}}(\neg\beta)$.
[…]
Let $rows(\beta)=\{v_{0},\ldots,v_{n}\}$ be the nmatrix of $\beta$ and let $rows(\gamma)=\{w_{0},\ldots,w_{n}\}$ be the nmatrix of $\gamma$. By (IH), we get the following
$lb_{\mathcal{M}}(\beta)\leq|rows(\beta)|\leq ub_{\mathcal{M}}(\gamma)$ (4)
$lb_{\mathcal{M}}(\gamma)\leq|rows(\gamma)|\leq ub_{\mathcal{M}}(\gamma)$ (5)
Using 4 and 5 and some laws of algebra, we get
$lb_{\mathcal{M}}(\beta)\times lb_{\mathcal{M}}(\gamma)\leq rows(\beta)\times rows(\gamma)\leq ub_{\mathcal{M}}(\beta)\times ub_{\mathcal{M}}(\gamma)$
Now, to construct the truth table for $\alpha$, we combine $rows(\beta)$ with $rows(\gamma)$. In the worst case, $sub(\beta)\cap sub(\gamma)=\emptyset$. In this case, we append, to each row of $rows(\beta)$, a row of $rows(\gamma)$. Then, each row creates $\kappa-1$ fresh rows. Therefore,
$|rows(\alpha)|\leq|rows(\beta)|\times|rows(\gamma)|\times\kappa$ (6)
Moreover,
$|sub(\alpha)|$ $\leq|sub(\beta)|+|sub(\gamma)|+1$ (7)
$|atoms(\alpha)|$ $\leq|atoms(\beta)|+|atoms(\gamma)|$ (8)
Therefore,
$|\mathcal{V}|^{|atoms(\alpha)|}$ $\leq|\mathcal{V}|^{|atoms(\beta)|+|atoms(\gamma)|}$
$=|\mathcal{V}|^{|atoms(\beta)|}\times|\mathcal{V}|^{|atoms(\gamma)|}$
$=lb_{\mathcal{M}}(\beta)\times lb_{\mathcal{M}}(\gamma)$
and
$\kappa^{(|sub(\alpha)|-|atoms(\alpha)|)}\times|\mathcal{V}|^{|atoms(\alpha)|}$
$\leq\kappa^{(|sub(\beta)|+|sub(\gamma)|+1)-(|atoms(\beta)|+|atoms(\gamma)|)}\times|\mathcal{V}|^{|atoms(\alpha)|}$
$=\kappa^{(|sub(\beta)|-|atoms(\beta)|)+(|sub(\gamma)|-|atoms(\gamma)|)+1}\times\left(|\mathcal{V}|^{|atoms(\beta)|}\times|\mathcal{V}|^{|atoms(\gamma)|}\right)$
$=\left(\kappa^{(|sub(\beta)|-|atoms(\beta)|)}\times|\mathcal{V}|^{|atoms(\beta)|}\right)\times\left(\kappa^{(|sub(\gamma)|-|atoms(\gamma)|)}\times|\mathcal{V}|^{|atoms(\gamma)|}\right)\times\kappa$
$=ub_{\mathcal{M}}(\beta)\times ub_{\mathcal{M}}(\gamma)\times\kappa$
Hence,
$lb_{\mathcal{M}}(\beta\#\gamma)\leq|rows(\beta\#\gamma)|\leq ub_{\mathcal{M}}(\beta\#\gamma)$
5.1 Algorithms
Given a formula $\varphi$, to decide if $\varphi$ is valid we have to proceed in two steps: (1) construct the initial matrix by applying the definitions of the multi-operations, and, then, (2) refine the matrix by removing every spurious row. In the previous section, we have proposed a way to measure the space complexity of (1). Now, we will find the time complexity of (2).
5.1.1 Pseudo-codes
In what follows, given an ordered pair $\langle a;b\rangle$,
$\pi_{1}(\langle a;b\rangle)=a$
and
$\pi_{2}(\langle a;b\rangle)=b$
Moreover, $\mathcal{L}$ denotes the set of every intuitionistic formula and $\mathcal{V}=\{\textbf{F},\textbf{U},\textbf{T}\}$.
Algorithm 1 Compatibility procedure. This algorithm receives two valuations, $v$ and $w$, and check if, for every evaluated formula $\alpha$, if $v(\alpha)=\textbf{T}$ then $w(\alpha)=\textbf{T}$ ($v$ is compatible with $w$). We assume that every valuation list is equally sorted.
procedure isCompatible($v,w:list(\mathcal{L}\times\mathcal{V})$)
     $i\leftarrow 0$
     while $i<length(v)$ do
         $a\leftarrow\pi_{2}(v[i])$triangleright Get the value in $v$ on index $i$.
         if $a=$ T then
              $b\leftarrow\pi_{2}(w[i])$triangleright Get the value in $w$ on the same index.
              if $b=$ T then
                  continue
              else
                  return $false$
              end if
         else
              continue
         end if
         $i\leftarrow i+1$
     end while
     return $true$
end procedure
Algorithm 2 Search algorithm. This algorithm receives a formula $A\in\mathcal{L}$, a (partial) valuation $v:\mathcal{L}\rightarrow\mathcal{V}$ and a matrix. It runs over the entire matrix looking for a valuation $w$ that validates $v$, that is, a $w$ such that $v$ is compatible with and such that $w(A)=\textbf{F}$.
procedure validate($A:\mathcal{L},v:list(\mathcal{L}\times\mathcal{V}),matrix:list(list(\mathcal{L}\times\mathcal{V}))$)
     $i\leftarrow 0$
     while $i<length(matrix)$ do
         $w\leftarrow matrix[i]$triangleright Get a valuation function.
         $j\leftarrow 0$
         while $j<length(w)$ do
              $A^{\prime}\leftarrow\pi_{1}(w[j])$triangleright Get a valuated formula.
              if $A^{\prime}=A$ then
                  $wA\leftarrow\pi_{2}(w[j])$triangleright Get $w(A)$.
                  if $wA=$ F then
                       $compatible\leftarrow isCompatible(v,w)$
                       if $compatible$ then
                           return $true$triangleright In this case, $w$ validates $v$.
                       else
                           continue
                       end if
                  else
                       continue
                  end if
              else
                  continue
              end if
              $j\leftarrow j+1$
         end while
         $i\leftarrow i+1$
     end while
     return $false$triangleright In this case, no $w$ validates $v$.
end procedure
Algorithm 3 Refinement algorithm. In this pseudo-code, we assume that $remove$ is a function that receives a matrix and a row, returning the matrix minus the row.
procedure refineMatrix($matrix:list(list(\mathcal{L}\times\mathcal{V}))$)
     $i\leftarrow 0$
     $newmatrix\leftarrow matrix$
     while $i<length(matrix)$ do
         $v\leftarrow matrix[i]$triangleright Get a valuation function.
         $j\leftarrow 0$
         while $j<length(v)$ do
              $A\leftarrow\pi_{1}(v[j])$triangleright Get a valuated formula.
              $vA\leftarrow\pi_{2}(v[j])$triangleright Get $v(A)$.
              if $vA=$ U then
                  $isValid\leftarrow validate(A,v,newmatrix)$
                  if $isValid$ then
                       continue $\triangleright$ In this case, we keep the matrix unaltered.
                  else
                       $newmatrix\leftarrow remove(v,newmatrix)$
                       break
                  end if
              else
                  continue
              end if
              $j\leftarrow j+1$
         end while
         $i\leftarrow i+1$
     end while
     return $newmatrix$triangleright Return the refined matrix.
end procedure
Algorithm 4 In the best case, the refinement algorithm does not remove any row. However, if it does drag at least one row, we have to rerun it, given that it could have been the case that the removed row was used to validate some other rows. For this reason, in the worst case, the refinement algorithm removes every row and one row at once. This auxiliar procedure should address this possibility.
procedure refineAux($matrix:list(list(\mathcal{L}\times\mathcal{V}))$)
     $i\leftarrow 0$
     $nrows\leftarrow length(matrix)$
     while $i<nrows$ do
         $matrix\leftarrow refineMatrix(matrix)$
         $i\leftarrow i+1$
     end while
     return $matrix$
end procedure
Theorem 5.6.
Given as input a matrix with $n$ rows and $m$ columns, the time complexity of refineAux is $O(n^{3}*m^{3})$.
Proof.
Given a list $v$ with length equal to $m$, in the worst case, the loop in $isCompatible$ runs $m$ times. Therefore, the complexity of this function is $O(m)$. Now, while, in the worst case, the external loop in $validate$ runs $n$ times, where $n$ is the number of rows in the matrix, the internal loop runs $m$ times. Because this functions calls $n*m$ times the function $isCompatible$, then the complexity of $validate$ is $O(n*m)*O(m)=O(n*m^{2})$. In the case of $refineMatrix$, the external loop runs $n$ times and the internal loop runs $m$ times. Therefore, it calls $validate$ $n*m$ times, which gives us $O(n*m)*O(n*m^{2})=O(n^{2}*m^{3})$. Finally, the function $refineAux$ calls $refineMatrix$ $n$ times. Therefore, the complexity of $refineAux$ is $O(n)*O(n^{2}*m^{3})=O(n^{3}*m^{3})$. ∎
5.2 Complexity of $\mathcal{R}(\mathcal{M}_{IPL})$
In the valuation of a formula in $\mathcal{M}_{IPL}$, we consider only $\mathbf{F}$ and $\mathbf{T}$ in the atomic cases. Given that, in this case, $\kappa=2$, we can simplify the upper bound to:
$ub_{\mathcal{M}_{IPL}}(\alpha)$ $=2^{(|sub(\alpha)|-|atoms(\alpha)|)}\times 2^{|atoms(\alpha)|}$ (9)
$=2^{|sub(\alpha)|}$ (10)
Now, given a formula $\varphi\in For(\Omega)$, assume that $|sub(\varphi)|=m$ and that $ub_{\mathcal{M}_{IPL}}(\varphi)=n$. Then
$n^{3}$ $=(2^{m})^{3}=8^{m}$
Therefore, given a formula $\varphi$, the time complexity to compute the truth table (in the worst case) in $\mathcal{R}(\mathcal{M}_{IPL})$ is given by $O(8^{m}*m^{3})$, which is in EXPTIME.
6 Related work and concluding remarks
In this paper, we proposed a new decision procedure for intuitionistic propositional logic IPL based on a 3-valued restricted Nmatrix semantics. Starting from the nice decision procedure for S4 obtained by Grätz in [4] based on a 3-valued RNmatrix, and taking advantage of the conservative translation of Gödel from IPL to S4, it is immediate to obtain an algorithm to decide validity in IPL. By means of an intermediate translation mapping from the language of IPL into the language of S4 we define a 3-valued Nmatrix which is the basis for an RNmatrix semantics for IPL. Moreover, by adapting Grätz’s algorithm for S4, a very simple algorithm for deleting spurious rows in the non-deterministic truth-table generated by the Nmatrix of IPL for a given formula is also provided. The sound and completeness proofs for the RNmatrix and for the obtained truth-tables algorithm for IPL, are presented with full technical details. These proofs were adapted from new ones for S4 we obtained in Subsection 3.2. On the one hand, we reduced the original Nmatrix for S4 presented by Grätz. On the other, we expanded it by adding the truth-tables for conjunction and disjunction. This required the extension and modification of the proofs originally given in [4]. It is worth noting that the Nmatrix for IPL we introduced in Definition 4.10 has a clear and intuitive interpretation in terms of constructibility, by observing the definition of the canonical valuation associated to a $\varphi$-saturated set in IPL given in Definition 4.19, as well as the observations made in Remark 4.20.
The idea of using finite non-deterministic truth-tables with restricted valuations for characterizing intuitionistic logic IPL is not new. Andréa Loparic already obtained in 1977, announced in [7] and finally published in [8], a semantical characterization of Johannson-Kolmogorov minimal logic and intuitionistic propositional logic IPL in terms of a $2$-valued, non truth-functional valuation semantics which induces a decision procedure for these logics. Similar to Grätz method (and to the adaptation to IPL proposed here), Loparic’s approach is based on a criterion for distinguish between that she calls semivaluations and valuations, the latter being the useful ones. This would correspond to level valuations, since they are exactly the semivaluations associate to $\varphi$-saturated sets, and the ones which characterizes the logic. The associated algorithm, as in the cases of S4 and IPL presented in this paper, generates a non-deterministic truth-table for a given formula in which, by using some formal criteria, some rows are deleted, and so the formula is tautology of the given logic if and only if it gets the value 1 in all the undeleted rows.
More recently, Alejandro Solares-Rojas proposed a 3-valued Nmatrix semantics for IPL based on intuitionistic Kripke models and the novel notion of depth-bounded approximations (see [11, Chapter 5]). Within this semantic approach to IPL, a 3-valued Nmatrix – whose truth-values are intuitionistic truth (1), falsity (0) and indeterminacy ($u$) – gives the meaning of the connectives in terms of information that a given agent holds. The valuations over the Nmatrix intend to model actual information states, which are ordered by a certain refinement relation. Such refinement relation aims to model the way knowledge of the agent evolves when new information, provided by reliable external sources, is acquired. The semantics for IPL is given by intuitionistic Kripke frames expanded by a family of valuations indexed by the worlds and where the refinement order corresponds to the accessibility relation of the frame. This interesting 3-valued non-deterministic semantics for IPL departs from a different perspective than ours. Besides the 3-valued Nmatrices proposed in [11, Chapter 5] and the one we give in Definition 4.10 are quite different, it is not immediate whether the semantics proposed by Solares-Rojas could be seen as given by an RNmatrix. This is the subject of future investigation.
It is interesting to notice that, by considering finite and decidable RNmatrix semantics (as the ones for S4 and IPL analyzed here) as a generalization of the standard notion of finite-valued (deterministic) truth-tables, the results presented here for S4 and IPL are a way to overcome the limiting results obtained by Gödel and Dugundji for IPL and S4, respectively, mentioned in Section 1.
As a topic of future research we intend to explore the meaning of the partial valuations corresponding to the rows of the non-deterministic truth-tables in IPL and S4. In particular, we plan to investigate their possible connection with the mini canonical models considered by Hughes and Cresswell in [5, Chapter 8] to prove the finite model property for some modal systems, including S4.
Finally, the extension of this methodology to other logics such as S5 and Prawitz’s ecumenical systems (see [10]) will also be investigated.
Acknowledgments: The first and second author thanks the support of the São Paulo Research Foundation (FAPESP, Brazil) through the PhD scholarship grant #21/01025-3 and the Thematic Project RatioLog #20/16353-3. The second author also acknowledges support from the National Council for Scientific and Technological Development (CNPq, Brazil), through the individual research grant #309830/2023-0.
Appendix A
The following truth-tables show that each axiom of the Hilbert calculus $\mathcal{H}_{IPL}$ for IPL is valid in the intuitionistic truth-tables. These tables were automatically generated by a Coq module for RNmatrices written by R. Leme, see
$q$ $p$ $q\to p$ $p\to(q\to p)$
F F U T
F F T T
F T T T
T F F T
T T T T
$p$ $q$ $p\land q$ $q\to p\land q$ $p\to(q\to p\land q)$
F F F U T
F F F T T
F T F F T
T F F T T
T T T T T
$q$ $p$ $r$ $q\to r$ $p\to q$ $p\to r$ $p\to(q\to r)$ $(p\to q)\to(p\to r)$ $(p\to(q\to r))\to((p\to q)\to(p\to r))$
F F F U U U U U T
F F F U U U T T T
F F F U U T T T T
F F F U T U U F T
F F F U T T T T T
F F F T U U T T T
F F F T U T T T T
F F F T T T T T T
F F T T U T T T T
F F T T T T T T T
F T F U F F F U T
F T F T F F T T T
F T T T F T T T T
T F F F T U U F T
T F F F T T T T T
T F T T T T T T T
T T F F T F F F T
T T T T T T T T T
$q$ $p$ $p\land q$ $p\land q\to p$ $p\land q\to q$
F F F T T
F T F T T
T F F T T
T T T T T
$p$ $q$ $p\lor q$ $p\to p\lor q$ $q\to p\lor q$
F F F T T
F T T T T
T F T T T
T T T T T
$p$ $q$ $r$ $p\to r$ $q\to r$ $p\lor q$ $p\lor q\to r$ $(q\to r)\to(p\lor q\to r)$ $(p\to r)\to((q\to r)\to((p\lor q)\to r))$
F F F U U F U U T
F F F U U F U T T
F F F U T F U F T
F F F T U F U T T
F F F T T F T T T
F F T T T F T T T
F T F U F T F T T
F T F T F T F T T
F T T T T T T T T
T F F F U T F U T
T F F F U T F T T
T F F F T T F F T
T F T T T T T T T
T T F F F T F T T
T T T T T T T T T
$p$ $q$ $q\to p$$\neg p$$\neg q$ $q\to\neg p$ $(q\to\neg p)\to\neg q$ $(q\to p)\to((q\to\neg p)\to\neg q)$
F F U U U U U T
F F U U U U T T
F F U U U T F T
F F U T U T F T
F F T U U U T T
F F T U T T T T
F F T T T T T T
F T F U F F U T
F T F U F F T T
F T F T F T F T
T F T F U U T T
T F T F T T T T
T T T F F F T T
$p$ $q$$\neg p$$\neg p\to q$ $p\to(\neg p\to q)$
F F U U T
F F U T T
F F T F T
F T U T T
F T T T T
T F F T T
T T F T T